### v1

# T-3DGS: Removing Transient Objects for 3D Scene Reconstruction

Abstract
We propose a novel framework to remove transient objects from input videos for 3D scene reconstruction using Gaussian Splatting. Our framework consists of the following steps. In the first step, we propose an unsupervised training strategy for a classification network to distinguish between transient objects and static scene parts based on their different training behavior inside the 3D Gaussian Splatting reconstruction. In the second step, we improve the boundary quality and stability of the detected transients by combining our results from the first step with an off-the-shelf segmentation method. We also propose a simple and effective strategy to track objects in the input video forward and backward in time. Our results show an improvement over the current state of the art in existing sparsely captured datasets and significant improvements in a newly proposed densely captured (video) dataset. More results and code are available at https://transient-3dgs.github.io.
Figure 1: Recent state-of-the-art methods, e.g. SpotLessSplats [34], often fail to correctly identify masks for transient / semi-transient objects. Our proposed T-3DGS method successfully locates all transient distractors, extracts clean masks, and propagates them between frames. By masking transient objects, T-3DGS can produce clean reconstruction results with high-fidelity novel view synthesis on a given sequence of images or videos captured in the wild.
1 Introduction
Novel view synthesis and 3D scene reconstruction from multiple 2D images or videos have emerged as rapidly evolving research areas within computer vision. Neural Radiance Fields (NeRF) [21] and, more recently, 3D Gaussian Splatting (3DGS) [12] have shown remarkable improvements in novel view synthesis on complex scenes. NeRF implicitly represents the scene as a volumetric function, and 3DGS explicitly represents it as a set of 3D Gaussians. Both approaches produce high-quality realistic images. There are multiple follow-up works for diverse downstream applications including 3D scene reconstruction [44, 16, 9], 3D synthesis [25, 48, 39], semantic and language integration into 3D representations [37, 13, 36], etc.
Both 3D Gaussian Splatting and NeRF optimize 3D scene reconstruction using photometric losses. High-quality results are achieved under the assumption that the captured scene is completely static and does not include any distractors, such as moving objects (i.e. transient objects), shadows, lightning changes, etc. However, in real-world scenarios, such an assumption can hardly be satisfied. The recordings, even when captured carefully, often contain moving people, cars, or other dynamic objects with their shadows, especially in locations that are tourist landmarks. Ignoring distractors during scene optimization leads to undesired blurring effects and floating artifacts. At the same time, removing such distractors from the captured recordings is challenging and limits NeRF and 3DGS widespread usage. The process of manual annotations of distractors is labor-intensive. Another approach is to utilize pre-trained segmentation models to locate transient distractors. This approach has two main limitations: 1) it needs prior knowledge of transients as a semantic class, and 2), more importantly, existing segmentation models cannot distinguish between static and dynamic objects of the same semantic class. Additionally, we would like to identify semi-transient objects in recordings and remove them from the scene. We define a semi-transient object as an object that has both dynamic and static states during the capturing process, e.g. a pushed chair stops after some time and becomes a fully static object. Therefore, we need more robust identification methods for transient and semi-transient distractors throughout the captured recordings.
In our work, we introduce a novel T-3DGS approach to 3D static scene reconstruction with 3DGS from a monocular video taken in uncontrolled setting. Our method comprises two main components: an unsupervised transient detector and a transient mask propagation framework. Firstly, as it was observed previously [31, 34], relying only on image residuals in transient identification is not a robust approach, e.g., due to appearance changes and color similarity between transient objects and background. We need more robust features, which are semantically consistent as well as novel architecture and loss design. Our proposed transient network is trained concurrently on both rendered (reconstructed) images and reference images through novel robust regularization losses. It utilizes semantically consistent features and is capable of producing more accurate masks with fewer misidentifications of transient objects, and significantly reduces the number of falsely identified objects as dynamic objects. Secondly, we have shown in our experiments that concurrent works [34, 43] fail to identify the semi-transient distractors and remove them from the scene as shown in Fig. 1. We further introduce our mask propagation framework that allows the extraction of object-aware masks that are consistent throughout the video and are especially helpful in improving the consistency of semi-transient distractor masks. Overall, our method remains robust for any type of distractors present in the scene. Lastly, we have collected a novel T-3DGS dataset containing several challenging scenes filled with semi-transient and slow-moving objects with dense frame rates. We evaluate our approach on casually captured scenes [31, 33] and on our semi-transient dataset. Our method consistently outperforms state-of-the-art approaches in reconstruction quality.
Our key contributions include:
  • A novel transient network to identify dynamic distractors utilizing features from a visual foundation model.
  • A novel video object segmentation module that robustly tracks objects throughout the video regardless of the frame rate and semi-transient behavior.
  • A novel challenging dataset that contains diverse scenes with semi-transient distractors and scenes with slow-moving objects for evaluation.
  • Our proposed method presents new state-of-the-art performance on robust reconstruction of the static scene parts on standard benchmark datasets.
2 Related Work
We briefly review Neural Radiance Fields and 3D Gaussian Splatting works with a focus on removing non-static distractors in the scene.
Neural Radiance Fields (NeRFs) [21] have emerged as a widely adopted method for high-quality scene reconstruction and novel view synthesis of 3D scenes. There are several follow-up extensions of NeRFs for fast training and inference [38, 22, 3], novel view synthesis on dynamic scenes [26, 42, 24], training on single or limited views [11, 30]. The seminal paper 3D Gaussian Splatting [12] employs Gaussian primitives to model scenes instead of relying on continuous volumetric representations. This method has recently gained popularity as a faster alternative to NeRFs. Follow-up works have extended 3DGS in many directions, for instance in dynamic [47, 52] and static [20, 50] scenes, utilizing additional information such as depth [55].
Handling Distractors in NeRFs. NeRF-W [19] and RobustNeRF [33] are two pioneering works similarly approaching the problem of handling transient distractors by utilizing photometric error. NeRF-W reconstructs both static background and transients combined with a data-dependent uncertainty field. RobustNeRF utilizes Iteratively Reweighted Least Squares for transient object identification and removal. Both methods rely directly on color residual supervision and often misclassify transient objects and backgrounds with similar colors. Additionally, they both require careful hyper-parameters tuning. Another alternative approach is NeRFBusters [46], which uses a data-driven prior to remove artifacts. NeRF On-the-go [31] utilizes DINOv2 features [23] to identify and eliminate distractors by predicting uncertainties through a shallow MLP and can deal with more complex scenes than RobustNeRF. NeRF-HuGS [2] utilizes two types of heuristics: 1) COLMAP-based [35] features combined with SAM [14] and 2) residual-based heuristics to identify and remove transient distractors. Their method is not robust to heavy transient distractions because both heuristics are not stable under such conditions, as shown in [31].
Extracting Features from Vision Foundation Models. Vision Foundation Models (VFMs) are trained on large-scale data to perform strong generalization ability to unseen domains or novel tasks. For example, there are a set of task-specific foundation models for image classification  [7, 41, 17], for segmentation [14, 29], and depth prediction [51] tasks. Task-agnostic models trained through self-distillation like DINO [1, 23] learn features that can be generalized for multiple vision tasks. CLIP [27] is a weakly-supervised multi-modal model that learns visual concepts from natural language supervision by contrastive pre-training on image-text pairs and performs zero-shot transfer to a wide range of downstream tasks. The features from Diffusion Models [10, 32] have also been used for semantic or keypoint correspondence [40, 54, 18]. In our work, we aim to utilize features from foundation models to identify transient distractors more robustly. We propose to use DINOv2 features [23] because they are semantically meaningful and spatially consistent, which is crucial for reliable distractor identification.
Video Object Segmentation. The goal of semi-supervised VOS is to identify when an object appears for the first time and then track it throughout the video. There are several recent approaches based on transformers [4, 5]. However, current methods suffer from mask inconsistencies throughout the video, especially when objects disappear from the video and reappear later. Additionally, current methods assume that the input is provided with a high frame rate, and these methods are not stable when the frame rate is low. In our work, we address these shortcomings.
Handling Distractors in 3DGS. There are multiple works addressing training 3DGS on unconstrained in-the-wild photo collections. SWAG [6] improves robustness of 3DGS by learning an appearance embedding space and image-dependent opacity variations to handle transient objects better. Gaussians in the Wild (GS-W) [53] utilizes CNN features to capture dynamic and intrinsic appearances from a reference image. Wild-GS [49] explicitly learns appearance embeddings by sampling the triplane from the reference image. Robust 3DGS [43] proposes a self-supervised approach to identify transient distractors by utilizing image residuals and leveraging a pre-trained segmentation network to produce object-aware masks. SpotLessSplats [34] proposes a method to identify transient objects by utilizing pre-computed feature maps from a foundation model [40] coupled with a robust optmization of 3DGS. These works [6, 53, 49, 43, 34] suffer from 1) hyper-parameter tuning, i.e., threshold parameters, 2) predicting inaccurate transient masks across the video and 3) relying on image residuals that result in falsely detecting transients as shown in Fig. 1. In our approach we try to address the limitations of the current works by identifying transients more accurately and consistently between video frames.
3 Method
Figure 2: The architecture of the proposed method. We use a modified version of 3D Gaussian Splatting (in green), with masked loss term $\mathcal{L}_{\text{masked}}$ specified in (Eq. 6). Every iteration, we begin by rendering a reconstruction of a randomly sampled reference image. Then we calculate RGB color residuals (i.e $L_{1}$ difference of those images), and DINOv2 features of both images. These features are then fed to our TMP model (in purple) that predicts binary masks for both reference and rendered images (shown in green on the actual images for clarity). We then calculate the loss specified in (Eq. 1) and update the parameters of the TMP via backpropagation, denoted with dashed lines. Notably, we detach rendered image masks to prevent the TMP from learning on features of the rendered image. Additionally, we use a SAM-based mask propagation module, as it further increases the consistency and sharpness of masks.
We propose a novel approach for reconstructing static scenes from unconstrained videos containing dynamic objects using 3D Gaussian Splatting (3DGS). Our method, illustrated in Fig. 2, introduces two key components to handle dynamic objects effectively: (1) transient mask predictor (TMP) and (2) transient mask refiner (TMR). The transient area detection component, implemented through our transient mask learning predictor, identifies regions containing dynamic objects by predicting per-pixel probabilities using semantic features, enhanced with a novel consistency regularization term. This regularization reduces false positive predictions while maintaining accurate detection of dynamic objects. The transient mask propagator component improves transient detection in both the spatial and temporal domain by leveraging SAM2 [29] to propagate transient masks across different frames, enabling artifact-free reconstruction.
3.1 3D Gaussian Splatting
We base our method on 3DGS [12]. Given a set of posed images $\{I_{n}\}_{n=1}^{N},I_{n}\in\mathbb{R}^{H\times W\times C}$, 3DGS represents a 3D scene as a set of anisotropic Gaussians $\{\mathcal{G}_{i}\}$, where each Gaussian is represented by its position (mean) $\mu_{i}$, a positive semi-definite covariance matrix $\Sigma_{i}$, an opacity $\alpha_{i}$, and a view-dependent appearance component (color) parametrized by spherical harmonics (SH) [28]. 3DGS representation is learned through optimization of Gaussian parameters via stochastic gradient descent.
Each 3D Gaussian is projected onto the image plane through a differentiable rasterization process to render an image from a specific viewpoint. First, the 3D Gaussian’s covariance matrix $\Sigma_{i}$ is projected to obtain a 2D covariance matrix $\Sigma^{\prime}_{i}$ in screen space: $\Sigma^{\prime}_{i}=JW\Sigma_{i}W^{T}J^{T}$, where $W$ is the perspective transformation matrix and $J$ is the Jacobian of the projection matrix. The projected 2D Gaussian’s contribution to each pixel $(x,y)$ is then computed as: $\alpha_{i}=\exp(-\frac{1}{2}(p-\mu^{\prime}_{i})^{T}(\Sigma^{\prime}_{i})^{-1}(p-\mu^{\prime}_{i}))$, where $p$ is the pixel coordinate and $\mu^{\prime}_{i}$ is the projected mean. The final color at each pixel is obtained through alpha compositing of all contributing Gaussians, sorted by depth: $C=\sum_{i=1}^{M}T_{i}\alpha_{i}c_{i}$, where $T_{i}=\prod_{j<i}(1-\alpha_{j})$ is the accumulated transmittance, $c_{i}$ is the view-dependent color computed from spherical harmonics coefficients, and $M$ is the number of Gaussians contributing to the pixel.
3.2 Transient Mask Prediction
Given the input images $\{I_{n}\}_{n=1}^{N}$, the goal is to optimize unsupervised TMP through 3DGS reconstruction to identify transient distractors without explicit supervision as shown in Fig. 2. The training process consists of two steps: (1) optimizing the 3DGS parameters while keeping the TMP frozen, and (2) optimizing the TMP while keeping the 3DGS parameters frozen. After each iteration, both models’ weights are updated. The TMP is trained to identify transient objects without explicit supervision, purely from the reconstruction objectives. Several recent works [6, 34, 15, 8, 49] follow a similar approach, which has demonstrated the effectiveness of such optimization for handling dynamic scenes.
Following [15, 34], we reformulate the transient detection problem as a semantic feature classification task rather than a direct pixel-level prediction. This approach leverages pre-trained foundation models to extract rich semantic features from images, enabling our system to make decisions based on high-level semantic understanding rather than solely relying on color information. This semantics-aware approach proves to be more robust in distinguishing between static and transient objects compared to traditional color-based methods. One of our main contributions is the unsupervised consistency loss in Eq. 3.
3.2.1 Feature Extraction
For each training iteration, we extract DINO v2 features [23] from both the input image $I$ and corresponding rendering $\hat{I}$ producing feature maps $f,\hat{f}$ respectively. We choose DINO v2 for several key reasons: (1) its self-supervised training enables robust semantic understanding without class-specific biases, (2) it demonstrates strong performance in distinguishing object boundaries and semantic regions even for previously unseen objects, (3) compared to alternatives like DIFT [40] features, DINO v2 offers significantly faster computation times, making it more practical for iterative training processes. These features provide a rich foundation for our TMP to make accurate decisions about scene dynamics without explicit supervision.
3.2.2 Transient Probability Masks
The extracted features are processed through our TMP module, which implements logistic regression to compute per-feature probability masks: $P^{f}=\sigma(\mathbf{W}\mathbf{f})$, where $\sigma(\cdot)$ is the sigmoid function, $\mathbf{W}$ represents the TMP weights. Since the features are patch-wise ($14\times 14$ px), we upsample predicted probability masks $P^{f}$ and $\hat{P}^{f}$ to the original image resolution using bilinear interpolation, obtaining pixel-wise transient masks $P$ and $\hat{P}$ respectively. Notably, TMP learns to distinguish transient objects without explicit supervision, guided only by the optimization objectives described below.
3.2.3 Loss Function
The TMP is optimized using three complementary losses:
$\mathcal{L}_{\text{transient}}=\mathcal{L}_{\text{RGB}}+\lambda_{prior}\mathcal{L}_{\text{reg}}+\mathcal{L}_{\text{consist}},$ (1)
The RGB loss identifies regions where the rendered image differs from the ground truth: $\mathcal{L}_{\text{RGB}}=\left|\left|(1-P)\odot(I-\hat{I})\right|\right|_{1}$. where $\|\cdot\|_{1}$ denotes the $L_{1}$ norm, and $\odot$ represents the element-wise (Hadamard) product. This loss exploits the observation that dynamic objects typically create larger discrepancies between rendered and ground truth images. By weighting these differences with $(1-P)$, we encourage the classifier to assign higher transient probabilities to regions showing significant reconstruction errors.
The regularization term promotes sparse predictions: $\mathcal{L}_{\text{reg}}=\left|\left|P\right|\right|_{1}$. This sparsity constraint helps produce compact, focused masks while reducing false positives.
Previous works  [8, 49] used a sum of $\mathcal{L}_{\text{RGB}}$ and $\lambda_{prior}\mathcal{L}_{\text{reg}}$ without the consistency loss, which has an explicit solution. For each pixel $(i,j)$, the optimal solution has the form:
$P^{*}_{i,j}=\begin{cases}1&\text{if }|I_{i,j}-\hat{I}_{i,j}|>\lambda_{prior}\\ 0&\text{if }|I_{i,j}-\hat{I}_{i,j}|\leq\lambda_{prior}\end{cases}$ (2)
However, this methodology exhibits significant limitations when reconstructing geometrically complex static structures. In such cases, even static objects may generate RGB errors exceeding $\lambda_{prior}$ due to reconstruction challenges, leading to their erroneous classification as dynamic elements and subsequent masking. This misclassification ultimately results in degraded reconstruction quality of the static scene components. To address these limitations, additional regularization constraints are necessary.
Our novel consistency term prevents misclassification in challenging regions:
$\mathcal{L}_{\text{consist}}=\left|\left|P\odot\hat{P}\right|\right|_{1}.$ (3)
This term addresses cases where static objects with complex surfaces might show high reconstruction error while being semantically consistent. The underlying principle is that while geometric complexity may lead to reconstruction errors in RGB space, the semantic features of static objects remain consistent between input and rendered images. By penalizing overlap between input and rendered image masks ($P$ and $\hat{P}$), we encourage the detector to focus on genuine semantic differences rather than reconstruction artifacts.
Fig. 3 demonstrates the effectiveness of this approach: after 7000 iterations, the model with consistency loss produces precise masks for transient objects in the input image while generating minimal masks for the rendered image. In contrast, without the consistency term, the TMP exhibits numerous false positives in both input and rendered images, often misclassifying poorly reconstructed static regions as transient objects. This behavior validates that the consistency loss effectively distinguishes between genuine dynamic objects and reconstruction artifacts based on their semantic representations.
Figure 3: The comparison of the TMP predictions with and without $\mathcal{L}_{\text{consist}}$ term. The quality of the predicted transient masks for the input image with consistency loss term is significantly better than without this term. The predicted masks by TMP are shown in green color.
3.2.4 Training Stability
To ensure stable training, we implement several key strategies. First, we detach $\hat{P}$ from the computational graph during consistency loss computation to prevent the model from learning to distinguish between ground truth and rendered images.
During 3DGS optimization, there are periods when renders may be unreliable: at the beginning of training and after each opacity reset. Following  [15], we address this by implementing two key strategies. First, we delay the start of TMP training until the 3DGS optimization has completed its first 500 iterations, ensuring the initial scene reconstruction has reached sufficient quality. Second, after each opacity reset, we temporarily pause the TMP optimization for 250 iterations while keep training 3DGS, allowing the reconstruction to stabilize before resuming transient detection.
Prior to bilinear interpolation, we dilate the masks by one pixel. This dilation step serves two purposes: it compensates for the coarse feature resolution of the foundation model (which processes images in patches) and helps address imprecise boundary predictions.
These modifications ensure robust training and accurate transient object detection across diverse dynamic scenes.
3.3 Transient Mask Refinement
Our transient area detection pipeline is robust for current benchmarks, where transient object is always dynamic and changes it’s position from frame to frame. But for semi-transient objects, which may do not change position for some frames it fails and masks only parts of video when it’s dynamic. To address these issue, we introduce a masks propagation process that refines transient masks into temporally consistent, accurate masks with high-resolution boundaries across the entire video sequence through refinement and propagation. Each segmentation consists of a binary mask defining the object’s spatial extent and a unique label that remains consistent throughout the video sequence.
3.3.1 Spatial Refinement
We employ the Segment Anything Model (SAM) [14] to refine our transient maps $P_{i}$ into more precise masks $M^{\prime}_{i}$. For each connected component $C_{i}^{k}$ in $P_{i}$, we sample up to ten points as prompts for SAM, leveraging its ability to generate high-quality segmentations from sparse inputs to extract a set of object-aware masks $M^{\prime}_{i}=\{M^{\prime}{{}_{i}^{j}}\}_{j=1}^{L_{i}}$, where $L_{i}$ is the number of predicted masks for image $I_{i}$. Due to potential inaccuracies in the boundaries of our masks, some sampled points might occasionally fall on the background rather than the object itself (e.g., a point sampled between the legs of a person). To address this issue, we filter the predicted masks based on their local coverage score:
$CS_{\text{local},i}=\frac{|P_{i}\cap M^{\prime}{{}_{i}^{j}}|}{|M^{\prime}{{}_{i}^{j}}|}$ (4)
We keep only masks that satisfy $CS_{\text{local},i}>\lambda_{\text{cov}}^{\text{ref}}$, forming the refined set $M^{\prime}_{i}=\{M^{\prime}{{}_{i}^{j}}|CS_{\text{local},i}>\lambda_{\text{cov}}^{\text{ref}}\}$.
3.3.2 Temporal Refinement
To address potential false negatives, we propagate the refined masks $\{M^{\prime}_{i}\}_{i=1}^{N}$ throughout the video using SAM2 [29] to obtain more consistent masks $\{M_{i}\}_{i=1}^{N}$. Our propagation process consists of three stages:
  1. 1.
    Forward Propagation: Iterating from the first to the last frame, to propagate segmentation masks forward.
  2. 2.
    Backward Propagation: Iterating Last-to-first frame, to propagate information from future frames backwards.
  3. 3.
    Final Propagation: Final Propagation: A final first-to-last pass, considering both past and future frames as context, which helps resolve any temporal inconsistencies.
To manage computational resources efficiently, we introduce a memory size parameter $N_{m}$, which limits the number of frames considered during propagation. At each step, we only maintain and use the segmentations from the $N_{m}$ most nearest frames, balancing between temporal consistency and memory constraints.
Throughout propagation, we manage mask intersections to ensure coherent segmentation. For any pair of masks $M_{i}^{l}$ and $M_{i}^{m}$ with $IoU(M_{i}^{l},M_{i}^{m})>\lambda_{merge}$, we merge them into a single mask, assigning the lower of the two original labels to maintain consistency.
3.3.3 Dynamic Object Filtration
To filter out false positive transients and ensure robust detection, we introduce a Stability Ratio (SR) metric that combines both spatial overlap accuracy and temporal consistency. For each detected object, the SR is calculated as $SR=\frac{1}{N}\sum_{i=1}^{N}(R_{i}\cdot CS_{\text{global},i})$, where $N$ is the number of valid frames, $R_{i}$ is the mean value of the absolute difference between ground truth and rendered images within the masked region in frame $i$, and $CS_{\text{global},i}=|P_{i}\cap M_{i}|/|M_{\text{max}}|$ is the global coverage score. Here, $P_{i}$ represents the prompt mask in frame $i$, $M_{i}$ is the segmentation mask, and $M_{\text{max}}$ is the maximum size of the object mask across all frames. This global score helps evaluate the object’s consistency relative to its largest observed size. A frame is considered valid and contributes to the SR calculation only if its local coverage score (Eq. 4) exceeds the validation threshold $\lambda_{\text{cov}}^{\text{val}}$. Objects with SR below a threshold $\lambda_{SR}$ are filtered out as potential false detections. This dual coverage score system ensures that objects maintain both spatial accuracy through local coverage and temporal consistency through global coverage and difference image values.
3.4 Artifact-Free Reconstruction
3DGS tends to generate floating artifacts (”floaters”) near the camera, particularly in challenging regions like those identified by transient masks. These artifacts can saturate gradients and degrade overall reconstruction quality. We address this issue through depth-aware regularization.
We render depth D for each pixel using alpha compositing similar to color rendering: $D=\sum_{i=1}^{M}T_{i}\alpha_{i}d_{i}$, where $d_{i}$ is the depth value of the i-th Gaussian, $T_{i}$ is the accumulated transmittance, and $\alpha_{i}$ is the opacity value. To suppress floating artifacts while preserving sharp depth discontinuities at object boundaries, we apply anisotropic total variation (TV) regularization to the rendered depth maps: $\mathcal{L}_{\text{depth}}=\text{mean}(|\nabla_{x}D|)+\text{mean}(|\nabla_{y}D|)$, where p indexes over pixels, $\nabla_{x}$ and $\nabla_{y}$ are spatial gradients in x and y directions respectively.
3.5 Masked Gaussian Splatting Optimization
The final step involves training the Gaussian Splatting model with the obtained masks $\{M_{i}\}_{i=1}^{N}$ for transients. Let $M_{i}$ be the binary mask for frame $i$, defined as:
$M_{i}(x,y)=\begin{cases}1&\text{if (x,y) is in an occluded area}\\ 0&\text{if (x,y) is in a static area},\end{cases}$ (5)
where (x,y) represents pixel coordinates in the image. We apply binary dilation to $M_{i}$ for $N_{e}$ iterations, yielding $M^{*}_{i}$. This operation creates a buffer zone around detected dynamic objects, enhancing the robustness of our static scene reconstruction. The final loss for 3DGS is:
$\begin{split}\mathcal{L}_{\text{masked}}=\lambda_{\text{SSIM}}\cdot L_{\text{SSIM}}(I_{i}\odot\overline{M}^{*}_{i},\hat{I_{i}}\odot\overline{M}^{*}_{i})+\\ \lambda_{\text{L1}}\cdot\left|\left|\overline{M}^{*}_{i}\odot(I_{i}-\hat{I_{i}})\right|\right|_{1}+\lambda_{\text{depth}}\mathcal{L}_{\text{depth}},\end{split}$ (6)
where $I_{i}$, $\hat{I_{i}}$, $\mathcal{L}_{depth}$, $\odot$ and $\|\cdot\|_{1}$ are the same as were defined previously, $L_{SSIM}$ is a structural similarity loss, $\overline{M}^{*}_{i}$ is a negation of $M^{*}_{i}$ that represents static background and $\lambda_{\text{SSIM}}$, $\lambda_{\text{L1}}$ and $\lambda_{\text{depth}}$ are weighting factors.
This formulation allows the model to focus on static scene elements, effectively handling dynamic objects in the reconstruction process. By integrating these steps, our method robustly reconstructs static scenes from unconstrained videos, effectively handling transient distractors.
4 Experiments
Mountain Fountain Corner Patio Spot Patio High Mean
PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
NeRF On-the-go [31] 20.15 0.64 20.11 0.61 24.22 0.81 20.78 0.75 23.33 0.79 21.41 0.72 21.67 0.72
3DGS [12] 20.09 0.70 20.11 0.69 23.51 0.85 15.87 0.68 20.78 0.80 17.80 0.69 19.69 0.74
Robust3DGS [43] 16.97 0.61 18.18 0.59 23.47 0.85 21.33 0.85 22.61 0.88 21.81 0.82 20.73 0.77
WildGaussians [15] 20.77 0.70 20.74 0.67 25.79 0.88 21.77 0.85 24.39 0.88 22.36 0.80 22.64 0.80
SpotLessSplats [34] 21.25 0.66 20.49 0.63 25.59 0.85 21.13 0.80 24.13 0.78 22.18 0.76 22.46 0.75
Ours 20.89 0.72 21.01 0.69 26.64 0.90 21.55 0.86 25.96 0.91 23.18 0.84 23.21 0.82
Table 1: Quantitative comparison on On-the-go dataset [31].
Lab 1 Lab 2 Library Anti-Stress Office Mean
PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
3DGS [12] 24.49 0.91 20.42 0.87 20.08 0.89 20.45 0.86 26.96 0.94 22.48 0.89
Robust3DGS [43] 25.35 0.93 24.74 0.92 24.33 0.93 22.95 0.91 28.52 0.96 25.18 0.93
WildGaussians [15] 25.71 0.92 23.68 0.91 24.65 0.92 21.69 0.89 28.89 0.95 24.92 0.92
SpotLessSplats [34] 25.28 0.91 24.63 0.90 24.11 0.91 22.22 0.90 28.08 0.92 24.86 0.91
Ours w/o TMR 25.69 0.93 24.74 0.92 23.82 0.92 24.21 0.92 29.01 0.95 25.49 0.93
Ours w/ TMR 27.71 0.95 25.42 0.93 28.34 0.97 28.79 0.97 29.87 0.96 28.03 0.96
Table 2: Quantitative comparison on our T-3DGS dataset.
We evaluate our proposed T-3DGS model on various datasets captured in uncontrolled settings and filled with diverse distractors. We perform qualitative and quantitative comparisons of our method with other SOTA methods. Finally, we provide an ablation study of our design and loss choices and discuss the limitations of the proposed method.
Datasets. We evaluate our model on three challenging datasets. The NeRF On-the-go dataset [31] contains 4 outdoor and 2 indoor sparsely captured scenes with different levels of occlusions (from 5% to over 30%) and minimal changes in appearance. The RobustNeRF dataset [33] contains 5 indoor scenes with unintentional changes during the capture process. Those changes include both transient objects that appear and disappear without forming a coherent video sequence (i.e., do not have a clear temporal ordering) and dynamic objects (e.g., floating balloons). Additionally, we introduce our novel T-3DGS dataset. The dataset contains 5 densely captured indoor scenes. Generally, dynamic objects in our videos are walking people and various small objects. However, unlike previous datasets, all scenes include various adversarial cases, such as transient, semi-transient, and slow-moving objects.
Baselines. We compare our model against vanilla 3D Gaussian Splatting [12] and against the current state-of-the-art method SpotLessSplats [34]. We further include WildGaussians [15] and Robust3DGS [43] as baselines. To compare different models we adopt commonly used PSNR and SSIM [45] metrics for evaluation.
Implementation details. All our experiments are conducted following the training setup from the official 3DGS implementation. We train our models for 30K iterations, using the Adam optimizer with a learning rate of 1e-3 for the TMP. The depth regularization loss $\mathcal{L}_{\text{depth}}$ is activated after the first 500 iterations, allowing the 3DGS to establish initial geometry reconstruction. For the experiments with mask propagation, we first train the TMP for 7000 iterations. At that point we pause the training to propagate the transient masks. Subsequently, we initiate a new training procedure using the propagated masks, keeping all other parameters the same as the original training setup. Thanks to our consistency term, we achieve robust results with a fixed regularization parameter $\lambda_{prior}=0.1$ across all scenes, eliminating the need for per-scene parameter tuning that was required in previous approaches.
Figure 4: Qualitative results on the On-the-go dataset. Our method detects transient objects better than our competitors. Predicted masks are shown in green.
Figure 5: Qualitative results on T-3DGS dataset. Our method detects cleaner transient masks and improves them further by mask propagation TMR module. Predicted masks are shown in green.
4.1 Quantitative Comparisons
We evaluate our model on all three datasets. We report results on On-the-go and T-3DGS datasets in Tab. 1 and 2, respectively, and we move evaluation results on RobustNeRF dataset to Supplementary Material as it is the least challenging. As shown in Tab. 1 and 2 our method generally outperforms current SOTA methods, in particular, our method is robust to changes in distant and high-frequency details. In Tab. 1 we run our method directly on masks predicted by TMP module without mask propagator.
While current SOTA methods struggle to detect semi-transient objects (Tab. 2), our proposed transient network TMP achieves higher performance minimizing false predictions. The integration of the SAM-based mask propagation TMR module further enhances our results in scene reconstruction filled with semi-transient objects, providing more accurate and reliable reconstructions.
4.2 Qualitative Comparisons
For qualitative comparison, we compare our method to SpotLessSplats [34] and Robust3DGS [43]. Fig. 4 and 5 show that our method has very few false negatives and robustly detects transients. For example, on the On-the-go dataset most of the methods struggle with high-frequency details and distant objects, as both generally reconstruct slower than the rest of the scene, and therefore confuse RGB residuals-based and similar methods. However, due to the introduction of our robust loss, such artifacts generally disappear from our dynamic maps. While our masks are less sharp, as we upsample them from low-resolution DINOv2 patches, techniques such as guided upsampling can improve performance of our method, however we leave this topic to following works. It should be noted, that SpotLessSplats [34] use features obtained from higher-resolution images, while we compute our features on a lower resolution, the same resolution used for training 3DGS.
For our T-3DGS dataset, we additionally utilize the SAM-based mask propagation module to propagate object-aware masks for semi-transient objects as shown in Fig. 5. While most methods would benefit from such a technique in principle, our masks have higher quality and fewer wrong detections. Applying mask propagation to other methods risks propagating errors.
On-the-go dataset
PSNR $\uparrow$ SSIM $\uparrow$
w/o dilation 22.65 0.82
w/o $\mathcal{L}_{\text{consist}}$ 16.76 0.69
w/o $\mathcal{L}_{\text{depth}}$ 23.02 0.82
Ours (full) 23.21 0.82
Table 3: We ablate the importance of each component of our method on On-the-go dataset. We report the mean over all scenes.
4.3 Ablation Study
We show the importance of each component in our T-3DGS model, separately disabling each of them. We report ablation results in Tab. 3 on On-the-go dataset. In the first experiment, we show the importance of the dilation before upsampling per-feature probability masks $P^{f}$. Without dilation the produced masks keep shrinking during the optimization process. This leads to a quality reduction in the renderings. In the second experiment, we turn off consistency loss term $\mathcal{L}_{\text{consist}}$. Our results show that adding this loss term improves the quality of renderings across all datasets. In our last experiment, we ablate the regularization term on depth $\mathcal{L}_{\text{depth}}$. Even though the metrics are only slightly improved, adding this term helps to remove floating artifacts in the reconstructed scene. We include visualizations to support our statements in Supplementary Material.
5 Conclusion
In this work, we have presented novel T-3DGS method for 3D scene reconstruction using Gaussian Splatting by effectively removing foreground dynamic distractors from input videos. The key to the success are unsupervised classification network to identify transient distractors through novel consistency loss term and the mask propagation approach that improves the boundary and temporal consistency of the predicted transient masks. Our results show that our method achieves significant improvement over the state-of-the-art methods on densely captured videos populated with semi-transient distractors. We believe our method is a valuable contribution towards widespread usage of 3DGS for robust 3D scene reconstruction from real-world videos captured in uncontrolled settings.
[…]
Supplementary Material
A Additional Implementation Details
For mask filtering and refinement in Sec. 3.3.1, we set $\lambda_{\text{cov}}^{\text{ref}}=0.7$ for initial mask refinement and $\lambda_{\text{cov}}^{\text{val}}=0.7$ for validation during object filtration. For temporal refinement in Sec. 3.3.2, we set the memory size parameter $N_{m}=10$ , which controls the number of frames considered during mask propagation. For the final mask dilation step, we perform $N_{e}=5$ iterations of binary dilation. Also, the mask merging threshold $\lambda_{merge}$ is set to 0.9, and the stability ratio threshold $\lambda_{SR}$ to 0.08 in Sec. 3.3.3.
B Evaluation on RobustNeRF Dataset
We provide evaluation on the RobustNeRF dataset [33]. As shown in Tab. 4 our method generally outperforms 3DGS [12], Robust 3DGS [43], WildGaussians [15], and shows similar performance comparing to SpotLessSplats [34]. We run our method directly on masks predicted by TMP module without mask propagator (TMR). In general, the dataset does not seem to be challenging enough to highlight the differences between different methods.
Android Statue Crab (1) Crab (2) Yoda Mean
PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
NeRF On-the-go [31] 23.50 0.75 21.58 0.77 - - - - 29.96 0.83 - -
3DGS [12] 23.51 0.81 21.35 0.84 30.39 0.94 31.53 0.92 29.80 0.92 27.32 0.89
Robust 3DGS [43] 24.40 0.83 22.10 0.85 34.41 0.96 32.99 0.93 32.62 0.93 29.30 0.90
WildGaussians [15] 24.89 0.83 22.69 0.87 30.16 0.93 31.11 0.91 30.50 0.91 27.87 0.89
SpotLessSplats [34] 24.45 0.79 22.50 0.80 35.45 0.95 33.29 0.94 33.55 0.94 29.85 0.88
Ours 25.04 0.84 22.90 0.87 35.81 0.95 33.36 0.93 32.09 0.92 29.84 0.90
Table 4: Quantitative comparison on RobustNeRF dataset [33].
Figure 6: Qualitative results on On-the-go dataset on the training frames. Our method produces higher-quality renderings without artifacts.
Figure 7: Qualitative results on On-the-go dataset on the testing frames. Our method produces higher-quality renderings without artifacts.
C More Qualitative Comparisons
For qualitative comparison, we compare our method to SpotLessSplats [34] and Robust3DGS [43]. We provide corresponding renderings for the masks shown in the main paper in Sec. 4.2. Fig. 6 and 7 show reconstructions of several scenes from the On-the-go dataset on training and testing frames, respectively. Currently, most methods can produce fairly good reconstructions and avoid significant artifacts, so generally, most methods produce fairly similar results (at least in the absence of semi-transient objects and other adversarial cases). However, it is notable that compared to other residual-based methods, we avoid misclassifying high-frequency details and similar objects. This improvement can be attributed to our consistency loss term.
Figure 8: We ablate the importance of each component of our method on On-the-go dataset on Corner scene. We provide qualitative comparison for masks (row 1), rendered images (row 2), and rendered depths (row 3).
Figure 9: We ablate the importance of each component of our method on On-the-go dataset on Spot scene. We provide qualitative comparison for masks (row 1), rendered images (row 2), and rendered depths (row 3).
D Additional Ablation Study
We provide visualizations that show the importance of each component in our proposed T-3DGS model on On-the-go dataset [31]. We show our ablation visualizations in Fig. 8 and 9.
Figure 10: The comparison of predicted masks for full and reduced scenes.
Figure 11: The comparison of predicted masks and scene reconstructions during movement of semi-transient objects for different frames.
E Handling Semi-Transient Objects
Semi-transient objects are have not been properly addressed in 3D scene reconstruction. Our method makes a large improvement over previous work and can handle fairly complex scenarios. We provide some details about data capture and the methodology we employed with respect to semi-transient objects. We also discuss why both our consistency term as well as mask propagation are important for handling semi-transient objects. Additionally, we discuss some limitations of our proposed method.
Some scenes in our dataset include reduced and full versions - in the reduced scenes the camera operator goes from one end of the scene to the other, while for the full scenes they also retrace their path and return to their initial positions, while semi-transient objects kept moving. As seen in Fig. 10, our proposed TMR module is crucial to produce good results even for the reduced scenes, the more challenging scenario. For full scenes, the additional data significantly improves mask predictions for all models. We conclude that handling semi-transient objects is an important challenge for in-the-wild video processing. In order to provide the most challenging datasets and to rigorously compare different methods - one must consider not only the kinds of motion of dynamic objects, but also their movement with respect to the camera.
As mentioned above, some of the scenes include semi-transient objects occluding the static scene for prolonged periods of time while remaining mostly still. As this period of time increases, semi-transient objects can effectively become static. While this effect might seem irrelevant to the detection of dynamic objects, this is not the case. As shown in the Fig. 11, most methods mask static background as if it was masking the semi-transient object. Notably, our method avoids this problem, and this highlights the importance of using both the consistency term and mask propagation algorithm we introduced. While the consistency term alone may still misclassify semi-transient objects, we avoid false classifications of static objects as dynamic. Therefore, our consistency term is crucial in order for the mask propagation to avoid introducing additional errors. This is in stark contrast to the competing methods, which have a lot more false positives. Mask propagation would introduce additional errors and may not help improving the overall quality.
Our method reliably removes transient and semi-transient distractors and successfully reconstructs static artifact-free 3D scenes. However, we have noticed that predicted masks are inflated due to the low resolution of the extracted feature maps, our method can give inconsistent results for small objects, since DINOv2 features are computed on patches. These problems could be addressed by utilizing feature extractors with higher-resolution feature maps or guided upsampling. Additionally, the temporal refinement process is constrained by a memory window of $N_{m}$ frames, which means that if an object disappears for more than $N_{m}$ frames and then reappears, it will be treated as a new instance with a different label. This can lead to inconsistent tracking and potentially affect the filtering process, especially for semi-transient objects that may temporarily leave the scene. Furthermore, our current filtering approach using global coverage scores may incorrectly filter out valid dynamic objects that undergo significant size changes, such as objects moving towards or away from the camera, or those experiencing perspective changes. We leave it as future work.

### v2

@@
 Abstract
-We propose a novel framework to remove transient objects from input videos for 3D scene reconstruction using Gaussian Splatting.
-Our framework consists of the following steps.
-In the first step, we propose an unsupervised training strategy for a classification network to distinguish between transient objects and static scene parts based on their different training behavior inside the 3D Gaussian Splatting reconstruction.
-In the second step, we improve the boundary quality and stability of the detected transients by combining our results from the first step with an off-the-shelf segmentation method.
-We also propose a simple and effective strategy to track objects in the input video forward and backward in time.
-Our results show an improvement over the current state of the art in existing sparsely captured datasets and significant improvements in a newly proposed densely captured (video) dataset.
-More results and code are available at https://transient-3dgs.github.io.
-Figure 1: Recent state-of-the-art methods, e.g.
-SpotLessSplats [34], often fail to correctly identify masks for transient / semi-transient objects.
-Our proposed T-3DGS method successfully locates all transient distractors, extracts clean masks, and propagates them between frames.
-By masking transient objects, T-3DGS can produce clean reconstruction results with high-fidelity novel view synthesis on a given sequence of images or videos captured in the wild.
+Transient objects in video sequences can significantly degrade the quality of 3D scene reconstructions.
+To address this challenge, we propose T-3DGS, a novel framework that robustly filters out transient distractors during 3D reconstruction using Gaussian Splatting.
+Our framework consists of two steps.
+First, we employ an unsupervised classification network that distinguishes transient objects from static scene elements by leveraging their distinct training dynamics within the reconstruction process.
+Second, we refine these initial detections by integrating an off-the-shelf segmentation method with a bidirectional tracking module, which together enhance boundary accuracy and temporal coherence.
+Evaluations on both sparsely and densely captured video datasets demonstrate that T-3DGS significantly outperforms state-of-the-art approaches, enabling high-fidelity 3D reconstructions in challenging, real-world scenarios.
+More results and code are available at https://transient-3dgs.github.io/
+Figure 1: Existing state-of-the-art methods, such as SpotLessSplats [22], often struggle to correctly identify transient and semi-transient objects, leading to artifacts in 3D scene reconstruction.
+Our proposed T-3DGS method accurately detects all transient distractors, generates clean masks, and propagates them across frames.
+By effectively masking transient objects, T-3DGS enables high-fidelity novel view synthesis and significantly improves reconstruction quality from real-world image and video sequences.
+22footnotetext: Correspondence to: a.komarichev@skoltech.ru.
 1 Introduction
-Novel view synthesis and 3D scene reconstruction from multiple 2D images or videos have emerged as rapidly evolving research areas within computer vision.
-Neural Radiance Fields (NeRF) [21] and, more recently, 3D Gaussian Splatting (3DGS) [12] have shown remarkable improvements in novel view synthesis on complex scenes.
+Novel view synthesis and 3D scene reconstruction from multiple 2D images or videos are critical, rapidly evolving areas in computer vision.
+Neural Radiance Fields (NeRF) [15] and 3D Gaussian Splatting (3DGS) [9] have shown remarkable improvements in novel view synthesis on complex scenes.
 NeRF implicitly represents the scene as a volumetric function, and 3DGS explicitly represents it as a set of 3D Gaussians.
 Both approaches produce high-quality realistic images.
-There are multiple follow-up works for diverse downstream applications including 3D scene reconstruction [44, 16, 9], 3D synthesis [25, 48, 39], semantic and language integration into 3D representations [37, 13, 36], etc.
+There are multiple follow-up works for diverse downstream applications, including 3D scene reconstruction [29, 13, 8], 3D synthesis [17, 31, 26], semantic and language integration into 3D representations [25, 10, 24].
 Both 3D Gaussian Splatting and NeRF optimize 3D scene reconstruction using photometric losses.
@@
 transient objects), shadows, lightning changes, etc.
-However, in real-world scenarios, such an assumption can hardly be satisfied.
-The recordings, even when captured carefully, often contain moving people, cars, or other dynamic objects with their shadows, especially in locations that are tourist landmarks.
-Ignoring distractors during scene optimization leads to undesired blurring effects and floating artifacts.
-At the same time, removing such distractors from the captured recordings is challenging and limits NeRF and 3DGS widespread usage.
-The process of manual annotations of distractors is labor-intensive.
+In real-world scenarios, this assumption can hardly be satisfied.
+Even when carefully captured, recordings often contain moving people, cars, or other dynamic objects with their shadows, especially in locations that are tourist landmarks.
+Ignoring distractors during scene optimization results in undesired blurring effects and floating artifacts.
+At the same time, removing such distractors from the captured recordings is challenging and limits the widespread usage of NeRF and 3DGS.
+Manually annotating distractors is labor intensive.
 Another approach is to utilize pre-trained segmentation models to locate transient distractors.
@@
 Therefore, we need more robust identification methods for transient and semi-transient distractors throughout the captured recordings.
-In our work, we introduce a novel T-3DGS approach to 3D static scene reconstruction with 3DGS from a monocular video taken in uncontrolled setting.
-Our method comprises two main components: an unsupervised transient detector and a transient mask propagation framework.
-Firstly, as it was observed previously [31, 34], relying only on image residuals in transient identification is not a robust approach, e.g., due to appearance changes and color similarity between transient objects and background.
-We need more robust features, which are semantically consistent as well as novel architecture and loss design.
-Our proposed transient network is trained concurrently on both rendered (reconstructed) images and reference images through novel robust regularization losses.
-It utilizes semantically consistent features and is capable of producing more accurate masks with fewer misidentifications of transient objects, and significantly reduces the number of falsely identified objects as dynamic objects.
-Secondly, we have shown in our experiments that concurrent works [34, 43] fail to identify the semi-transient distractors and remove them from the scene as shown in Fig.
-1.
-We further introduce our mask propagation framework that allows the extraction of object-aware masks that are consistent throughout the video and are especially helpful in improving the consistency of semi-transient distractor masks.
-Overall, our method remains robust for any type of distractors present in the scene.
-Lastly, we have collected a novel T-3DGS dataset containing several challenging scenes filled with semi-transient and slow-moving objects with dense frame rates.
-We evaluate our approach on casually captured scenes [31, 33] and on our semi-transient dataset.
-Our method consistently outperforms state-of-the-art approaches in reconstruction quality.
-Our key contributions include:
-  • A novel transient network to identify dynamic distractors utilizing features from a visual foundation model.
-  • A novel video object segmentation module that robustly tracks objects throughout the video regardless of the frame rate and semi-transient behavior.
-  • A novel challenging dataset that contains diverse scenes with semi-transient distractors and scenes with slow-moving objects for evaluation.
-  • Our proposed method presents new state-of-the-art performance on robust reconstruction of the static scene parts on standard benchmark datasets.
+We introduce T-3DGS, a novel approach for 3D static scene reconstruction from monocular video in uncontrolled settings.
+Our method includes an unsupervised transient detector and a transient mask propagation framework.
+Relying solely on image residuals for transient identification is unreliable due to issues such as appearance changes and color similarity to the background [20, 22].
+To address this issue, we develop a divergence-based technique on top of the uncertainty modeling approach [12] to detect transients.
+It helps improve mask accuracy and significantly reduce misclassifications of transient objects.
+Our experiments show that concurrent works [22, 28] fail to remove semi-transient distractors (Fig.
+1).
+We introduce a mask propagation framework for extracting object-aware masks that improve consistency in case of semi-transient distractors.
+Our method remains robust to all types of distractors.
+Additionally, we present the novel T-3DGS dataset with challenging scenes featuring semi-transient and slow-moving objects.
+Evaluations on both casual scenes [20, 21] and our dataset show our method outperforms state-of-the-art approaches in reconstruction quality.
+Our key contributions, which together ensure consistent detection and removal of transient objects for improved 3D reconstruction, include:
+  • Generalized uncertainty modeling for efficient transient object identification;
+  • A divergence-based approach that leverages semantic consistency between reference and reconstructed frames for identifying transient objects;
+  • A robust video object segmentation module that tracks objects across varying frame rates and semi-transient behaviors;
+  • A challenging new dataset featuring diverse scenes with semi-transient distractors and slow-moving objects;
+  • State-of-the-art performance on benchmark datasets for robust static scene reconstruction.
 2 Related Work
-We briefly review Neural Radiance Fields and 3D Gaussian Splatting works with a focus on removing non-static distractors in the scene.
-Neural Radiance Fields (NeRFs) [21] have emerged as a widely adopted method for high-quality scene reconstruction and novel view synthesis of 3D scenes.
-There are several follow-up extensions of NeRFs for fast training and inference [38, 22, 3], novel view synthesis on dynamic scenes [26, 42, 24], training on single or limited views [11, 30].
-The seminal paper 3D Gaussian Splatting [12] employs Gaussian primitives to model scenes instead of relying on continuous volumetric representations.
+We provide a brief review of the works on Neural Radiance Fields and 3D Gaussian Splatting with a focus on removing non-static distractors in the scene.
+Neural Radiance Fields (NeRFs) [15] are widely adopted methods for high-quality scene reconstruction and novel view synthesis of 3D scenes.
+The seminal work 3D Gaussian Splatting [9] employs Gaussian primitives to model scenes instead of relying on continuous volumetric representations.
 This method has recently gained popularity as a faster alternative to NeRFs.
-Follow-up works have extended 3DGS in many directions, for instance in dynamic [47, 52] and static [20, 50] scenes, utilizing additional information such as depth [55].
 Handling Distractors in NeRFs.
-NeRF-W [19] and RobustNeRF [33] are two pioneering works similarly approaching the problem of handling transient distractors by utilizing photometric error.
+NeRF-W [14] and RobustNeRF [21] are two pioneering works approaching the problem in a similar way.
 NeRF-W reconstructs both static background and transients combined with a data-dependent uncertainty field.
 RobustNeRF utilizes Iteratively Reweighted Least Squares for transient object identification and removal.
-Both methods rely directly on color residual supervision and often misclassify transient objects and backgrounds with similar colors.
+Both methods rely on color residual supervision and frequently misclassify transient objects and backgrounds that share similar colors.
 Additionally, they both require careful hyper-parameters tuning.
-Another alternative approach is NeRFBusters [46], which uses a data-driven prior to remove artifacts.
 NeRF On-the-go [20] utilizes DINOv2 features [16] to identify and eliminate distractors by predicting uncertainties through a shallow MLP and can deal with more complex scenes than RobustNeRF.
 NeRF-HuGS [2] utilizes two types of heuristics: 1) COLMAP-based [23] features combined with SAM [11] and 2) residual-based heuristics to identify and remove transient distractors.
-Their method is not robust to heavy transient distractions because both heuristics are not stable under such conditions, as shown in [31].
+Their method lacks robustness to heavy transient distractions, as both heuristics are unstable under such conditions, as demonstrated in [20].
 Extracting Features from Vision Foundation Models.
-Vision Foundation Models (VFMs) are trained on large-scale data to perform strong generalization ability to unseen domains or novel tasks.
-For example, there are a set of task-specific foundation models for image classification  [7, 41, 17], for segmentation [14, 29], and depth prediction [51] tasks.
+Vision Foundation Models (VFMs) are trained on large-scale data, enabling strong generalization to unseen domains or novel tasks.
 Task-agnostic models trained through self-distillation like DINO [1, 16] learn features that can be generalized for multiple vision tasks.
-CLIP [27] is a weakly-supervised multi-modal model that learns visual concepts from natural language supervision by contrastive pre-training on image-text pairs and performs zero-shot transfer to a wide range of downstream tasks.
-The features from Diffusion Models [10, 32] have also been used for semantic or keypoint correspondence [40, 54, 18].
-In our work, we aim to utilize features from foundation models to identify transient distractors more robustly.
-We propose to use DINOv2 features [23] because they are semantically meaningful and spatially consistent, which is crucial for reliable distractor identification.
 Video Object Segmentation.
 The goal of semi-supervised VOS is to identify when an object appears for the first time and then track it throughout the video.
-There are several recent approaches based on transformers [4, 5].
-However, current methods suffer from mask inconsistencies throughout the video, especially when objects disappear from the video and reappear later.
-Additionally, current methods assume that the input is provided with a high frame rate, and these methods are not stable when the frame rate is low.
+Several recent approaches based on transformers [3, 4] have been proposed.
+However, current methods suffer from mask inconsistencies, particularly when objects disappear and reappear in the video.
+Additionally, these methods assume that the input has a high frame rate, and they become unstable when the frame rate is low.
 In our work, we address these shortcomings.
+Figure 2: Overview of the Proposed T-3DGS Architecture.
+We introduce a modified version of 3D Gaussian Splatting, incorporating a masked loss term $\mathcal{L}_{\text{masked}}$ as described in Eq.
+13.
+In each iteration, we start by rendering a reconstruction of a randomly sampled reference image.
+We compute residuals, along with DINOv2 features from both the ground truth and rendered images.
+These features are then fed to our RUP model to predict the per-pixel covariance matrix for both images.
+We calculate binary masks based on the divergence of these distributions (as specified in Eq.
+10).
+Subsequently, we compute the likelihood as described in Eq.
+7 and update the parameters of the RUP model via backpropagation, as indicated by the dashed lines.
+Additionaly, for some scenes, we incorporate a SAM-based mask refiner module (TMR), which further enhances the consistency and sharpness of the masks.
 Handling Distractors in 3DGS.
-There are multiple works addressing training 3DGS on unconstrained in-the-wild photo collections.
+Several works address the training of 3DGS on unconstrained, in-the-wild photo collections.
 SWAG [5] improves robustness of 3DGS by learning an appearance embedding space and image-dependent opacity variations to handle transient objects better.
@@
 SpotLessSplats [22] proposes a method to identify transient objects by utilizing pre-computed feature maps from a foundation model [27] coupled with a robust optmization of 3DGS.
-These works [6, 53, 49, 43, 34] suffer from 1) hyper-parameter tuning, i.e., threshold parameters, 2) predicting inaccurate transient masks across the video and 3) relying on image residuals that result in falsely detecting transients as shown in Fig.
+These works [5, 33, 32, 28, 22] suffer from: 1) the need for hyper-parameter tuning, such as threshold parameters; 2) inaccurate prediction of transient masks across the video; and 3) reliance on image residuals, leading to the false detection of transients, as shown in Fig.
 1.
-In our approach we try to address the limitations of the current works by identifying transients more accurately and consistently between video frames.
+In our approach, we aim to address the limitations of the current works by identifying transients more accurately and consistently across video frames.
 3 Method
-Figure 2: The architecture of the proposed method.
-We use a modified version of 3D Gaussian Splatting (in green), with masked loss term $\mathcal{L}_{\text{masked}}$ specified in (Eq.
-6).
-Every iteration, we begin by rendering a reconstruction of a randomly sampled reference image.
-Then we calculate RGB color residuals (i.e $L_{1}$ difference of those images), and DINOv2 features of both images.
-These features are then fed to our TMP model (in purple) that predicts binary masks for both reference and rendered images (shown in green on the actual images for clarity).
-We then calculate the loss specified in (Eq.
-1) and update the parameters of the TMP via backpropagation, denoted with dashed lines.
-Notably, we detach rendered image masks to prevent the TMP from learning on features of the rendered image.
-Additionally, we use a SAM-based mask propagation module, as it further increases the consistency and sharpness of masks.
-We propose a novel approach for reconstructing static scenes from unconstrained videos containing dynamic objects using 3D Gaussian Splatting (3DGS).
+We propose a novel approach to reconstructing static scenes from unconstrained videos that contain dynamic objects, utilizing 3D Gaussian Splatting (3DGS).
 Our method, illustrated in Fig.
-2, introduces two key components to handle dynamic objects effectively: (1) transient mask predictor (TMP) and (2) transient mask refiner (TMR).
-The transient area detection component, implemented through our transient mask learning predictor, identifies regions containing dynamic objects by predicting per-pixel probabilities using semantic features, enhanced with a novel consistency regularization term.
-This regularization reduces false positive predictions while maintaining accurate detection of dynamic objects.
-The transient mask propagator component improves transient detection in both the spatial and temporal domain by leveraging SAM2 [29] to propagate transient masks across different frames, enabling artifact-free reconstruction.
+2, introduces two key components designed to handle dynamic objects effectively: (1) reconstruction uncertainty predictor (RUP), and (2) transient mask refiner (TMR).
+The transient area detection component, implemented through our transient mask learning predictor, identifies regions containing dynamic objects by predicting per-pixel probabilities using semantic features.
+The transient mask refiner improves transient detections in both spatial and temporal domains by leveraging SAM2 [19] to propagate transient masks across multiple frames, facilitating artifact-free reconstruction.
 3.1 3D Gaussian Splatting
-We base our method on 3DGS [12].
+Our method is based on 3DGS [9].
 Given a set of posed images $\{I_{n}\}_{n=1}^{N},I_{n}\in\mathbb{R}^{H\times W\times C}$, 3DGS represents a 3D scene as a set of anisotropic Gaussians $\{\mathcal{G}_{i}\}$, where each Gaussian is represented by its position (mean) $\mu_{i}$, a positive semi-definite covariance matrix $\Sigma_{i}$, an opacity $\alpha_{i}$, and a view-dependent appearance component (color) parametrized by spherical harmonics (SH) [18].
@@
 First, the 3D Gaussian’s covariance matrix $\Sigma_{i}$ is projected to obtain a 2D covariance matrix $\Sigma^{\prime}_{i}$ in screen space: $\Sigma^{\prime}_{i}=JW\Sigma_{i}W^{T}J^{T}$, where $W$ is the perspective transformation matrix and $J$ is the Jacobian of the projection matrix.
-The projected 2D Gaussian’s contribution to each pixel $(x,y)$ is then computed as: $\alpha_{i}=\exp(-\frac{1}{2}(p-\mu^{\prime}_{i})^{T}(\Sigma^{\prime}_{i})^{-1}(p-\mu^{\prime}_{i}))$, where $p$ is the pixel coordinate and $\mu^{\prime}_{i}$ is the projected mean.
-The final color at each pixel is obtained through alpha compositing of all contributing Gaussians, sorted by depth: $C=\sum_{i=1}^{M}T_{i}\alpha_{i}c_{i}$, where $T_{i}=\prod_{j<i}(1-\alpha_{j})$ is the accumulated transmittance, $c_{i}$ is the view-dependent color computed from spherical harmonics coefficients, and $M$ is the number of Gaussians contributing to the pixel.
-3.2 Transient Mask Prediction
-Given the input images $\{I_{n}\}_{n=1}^{N}$, the goal is to optimize unsupervised TMP through 3DGS reconstruction to identify transient distractors without explicit supervision as shown in Fig.
+The contribution of projected 2D Gaussian to each pixel $(x,y)$ is computed as: $\alpha_{i}=\exp(-\frac{1}{2}(p-\mu^{\prime}_{i})^{T}(\Sigma^{\prime}_{i})^{-1}(p-\mu^{\prime}_{i}))$, where $p$ is the pixel coordinate and $\mu^{\prime}_{i}$ is the projected mean.
+The final color at each pixel is obtained by alpha compositing the contributions from all Gaussians, sorted by depth: $C=\sum_{i=1}^{M}T_{i}\alpha_{i}c_{i}$, where $T_{i}=\prod_{j<i}(1-\alpha_{j})$ is the accumulated transmittance, $c_{i}$ is the view-dependent color computed from spherical harmonics coefficients, and $M$ is the number of Gaussians contributing to the pixel.
+3.2 Reconstruction Uncertainty Prediction
+Given the input images $\{I_{n}\}_{n=1}^{N}$, the goal is to optimize the unsupervised reconstruction uncertainty predictor RUP through 3DGS reconstruction to identify transient distractors without explicit supervision as shown in Fig.
 2.
-The training process consists of two steps: (1) optimizing the 3DGS parameters while keeping the TMP frozen, and (2) optimizing the TMP while keeping the 3DGS parameters frozen.
-After each iteration, both models’ weights are updated.
-The TMP is trained to identify transient objects without explicit supervision, purely from the reconstruction objectives.
-Several recent works [6, 34, 15, 8, 49] follow a similar approach, which has demonstrated the effectiveness of such optimization for handling dynamic scenes.
-Following [15, 34], we reformulate the transient detection problem as a semantic feature classification task rather than a direct pixel-level prediction.
-This approach leverages pre-trained foundation models to extract rich semantic features from images, enabling our system to make decisions based on high-level semantic understanding rather than solely relying on color information.
-This semantics-aware approach proves to be more robust in distinguishing between static and transient objects compared to traditional color-based methods.
-One of our main contributions is the unsupervised consistency loss in Eq.
-3.
+Following prior research [14, 20, 12] we employ uncertainty modeling techniques, with significant modifications.
+RUP is trained to identify transient objects without explicit supervision, purely from the reconstruction objectives.
+Several recent works [5, 22, 12, 7, 32] follow a similar approach, demonstrating the effectiveness of this optimization in handling dynamic scenes.
+As in WildGaussians [12] (and, in contrast to NeRF counterparts [14, 20]) every iteration we update both Gaussian Splatting and RUP weights.
+Additionally, we detach masks when updating Gaussian Splatting and detach reconstructed images when updating RUP.
+Following [12, 22], we reformulate the transient detection problem as a semantic feature classification task.
+This approach leverages pre-trained foundation models to extract rich semantic features from images.
+By doing so, it enables our system to make decisions based on high-level semantic understanding, rather than relying solely on color information.
+This semantics-aware approach is more robust in distinguishing between static and transient objects than traditional color-based methods.
 3.2.1 Feature Extraction
-For each training iteration, we extract DINO v2 features [23] from both the input image $I$ and corresponding rendering $\hat{I}$ producing feature maps $f,\hat{f}$ respectively.
+For each training iteration, we extract DINOv2 features [16] from both the input image $I$ and the corresponding rendering $\hat{I}$, producing feature maps $f,\hat{f}$ respectively.
 We choose DINOv2 for several key reasons: (1) its self-supervised training enables robust semantic understanding without class-specific biases, (2) it demonstrates strong performance in distinguishing object boundaries and semantic regions even for previously unseen objects, (3) compared to alternatives like DIFT [27] features, DINOv2 offers significantly faster computation times, making it more practical for iterative training processes.
-These features provide a rich foundation for our TMP to make accurate decisions about scene dynamics without explicit supervision.
-3.2.2 Transient Probability Masks
-The extracted features are processed through our TMP module, which implements logistic regression to compute per-feature probability masks: $P^{f}=\sigma(\mathbf{W}\mathbf{f})$, where $\sigma(\cdot)$ is the sigmoid function, $\mathbf{W}$ represents the TMP weights.
-Since the features are patch-wise ($14\times 14$ px), we upsample predicted probability masks $P^{f}$ and $\hat{P}^{f}$ to the original image resolution using bilinear interpolation, obtaining pixel-wise transient masks $P$ and $\hat{P}$ respectively.
-Notably, TMP learns to distinguish transient objects without explicit supervision, guided only by the optimization objectives described below.
-3.2.3 Loss Function
-The TMP is optimized using three complementary losses:
-$\mathcal{L}_{\text{transient}}=\mathcal{L}_{\text{RGB}}+\lambda_{prior}\mathcal{L}_{\text{reg}}+\mathcal{L}_{\text{consist}},$ (1)
-The RGB loss identifies regions where the rendered image differs from the ground truth: $\mathcal{L}_{\text{RGB}}=\left|\left|(1-P)\odot(I-\hat{I})\right|\right|_{1}$.
-where $\|\cdot\|_{1}$ denotes the $L_{1}$ norm, and $\odot$ represents the element-wise (Hadamard) product.
-This loss exploits the observation that dynamic objects typically create larger discrepancies between rendered and ground truth images.
-By weighting these differences with $(1-P)$, we encourage the classifier to assign higher transient probabilities to regions showing significant reconstruction errors.
-The regularization term promotes sparse predictions: $\mathcal{L}_{\text{reg}}=\left|\left|P\right|\right|_{1}$.
-This sparsity constraint helps produce compact, focused masks while reducing false positives.
-Previous works  [8, 49] used a sum of $\mathcal{L}_{\text{RGB}}$ and $\lambda_{prior}\mathcal{L}_{\text{reg}}$ without the consistency loss, which has an explicit solution.
-For each pixel $(i,j)$, the optimal solution has the form:
-$P^{*}_{i,j}=\begin{cases}1&\text{if }|I_{i,j}-\hat{I}_{i,j}|>\lambda_{prior}\\ 0&\text{if }|I_{i,j}-\hat{I}_{i,j}|\leq\lambda_{prior}\end{cases}$ (2)
-However, this methodology exhibits significant limitations when reconstructing geometrically complex static structures.
-In such cases, even static objects may generate RGB errors exceeding $\lambda_{prior}$ due to reconstruction challenges, leading to their erroneous classification as dynamic elements and subsequent masking.
-This misclassification ultimately results in degraded reconstruction quality of the static scene components.
-To address these limitations, additional regularization constraints are necessary.
-Our novel consistency term prevents misclassification in challenging regions:
-$\mathcal{L}_{\text{consist}}=\left|\left|P\odot\hat{P}\right|\right|_{1}.$ (3)
-This term addresses cases where static objects with complex surfaces might show high reconstruction error while being semantically consistent.
-The underlying principle is that while geometric complexity may lead to reconstruction errors in RGB space, the semantic features of static objects remain consistent between input and rendered images.
-By penalizing overlap between input and rendered image masks ($P$ and $\hat{P}$), we encourage the detector to focus on genuine semantic differences rather than reconstruction artifacts.
+These features serve as a robust foundation, enabling RUP to make accurate decisions about scene dynamics without explicit supervision.
+3.2.2 Transient 2D Uncertainty Modeling
+As previously discussed, most methods detect transient objects by utilizing reconstruction errors. For example, NeRF On-the-go [20] considers RGB residuals:
+$R=||\hat{I}-I||_{2}.$ (1)
+It assumes that the residuals follow a normal distribution:
+$p(R|\sigma)=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{R^{2}}{2\sigma^{2}}\right).$ (2)
+Therefore, we can obtain negative log likelihood:
+$\mathcal{L}_{u}=\frac{R^{2}}{2\sigma^{2}}+\log\sigma+\frac{\log 2\pi}{2}.$ (3)
+Although the approach is reasonable, RGB residuals lack robustness.
+In particular, high-frequency objects often result in high reconstruction errors, producing incorrect misclassification.
+Similarly, dynamic objects with colors similar to the background may be classified as static.
+While DSSIM or DINOv2 cosine distance can mitigate some errors, they introduce their own limitations.
+DSSIM residuals are susceptible to similar errors as RGB residuals, though to a lesser degree.
+The DINOv2 cosine distance, while highly robust, suffers from low resolution.
+Upsampling models, such as FeatUP [6], can address this issue, though they introduce upsampling artifacts.
+This motivates a new multivariate formulation of uncertainty modeling. Let the residual be a 2-dimensional vector:
+$R=\begin{bmatrix}R_{1}\\ R_{2}\end{bmatrix},$ (4)
+where $R_{1}$ and $R_{2}$ correspond to different similarity metrics: 1) DINOv2 cosine distance defined like in WildGaussians, except we upscale it with FeatUP [6] and 2) DSSIM.
+We consider a multivariate normal distribution with zero mean and covariance matrix $\Sigma$:
+$p(R)=\frac{1}{(2\pi)\sqrt{|\Sigma|}}\exp\left(-\frac{1}{2}R^{T}\Sigma^{-1}R\right),$ (5)
+where the covariance matrix is given by
+$\Sigma=\begin{bmatrix}\sigma_{1}^{2}&\rho\sigma_{1}\sigma_{2}\\ \rho\sigma_{1}\sigma_{2}&\sigma_{2}^{2}\end{bmatrix}.$ (6)
+The negative log-likelihood function becomes:
+$\mathcal{L}_{u}=-\log p(R)=\frac{1}{2}R^{T}\Sigma^{-1}R+\frac{1}{2}\log|\Sigma|+\log 2\pi.$ (7)
+In contrast to previous works  [14, 20, 12] we predict three parameters — $\sigma_{1}$, $\sigma_{2}$, $\rho$ instead of a single $\sigma$. This allows us to combine information about both residuals.
+There is a problem in this derivation due to the assumption that residuals to be strictly positive.
+This implies that our distribution represents only the positive quadrant of the bivariate normal distribution.
+While this is relevant for the subsequent derivations, it does not affect the likelihood, as it only introduces a scaling factor, which can be ignored during optimization.
+Interestingly, this is not important for the one-dimensional case, where likelihood depends only on the absolute value of the residual.
+We train a neural network that takes DINOv2 features from a reference image as input and makes a per pixel prediction of $\Sigma$, and use our likelihood term in Eq.
+7 as a loss function.
+It should be noted that $\Sigma$ can be noninvertible when $\sigma_{i}=0$ or $\rho=\pm 1$.
+Although we predict $\sigma_{i}$ using a softplus nonlinearity, and $\rho$ using a tanh nonlinearity to avoid undesirable values, in practice this can lead to 1) numerical instabilities and 2) undesirable values due to the discrete representation of numerical values.
+The second problem is easy to solve with clamping, in our experience, the first problem was mostly solved by introducing normalization layers into the architecture of the neural network.
+3.2.3 Model Architecture
+Given that our training objective is considerably more challenging than the one-dimensional modeling of WildGaussians [12], our model requires a larger architecture.
+However, this also offers an advantage over previous methods, as we can use simple upscale layers to make our prediction denser without sacrificing local/nonlocal smoothing.
+The details of architecture are provided in the Supplementary Material.
+3.2.4 Binary Mask
+One approach to obtaining a binary mask using the modeled uncertainty is to set a threshold on one of the predicted values or define a new criterion:
+$M=\mathbb{I}(f(\sigma_{1},\sigma_{2},\rho)>C),$ (8)
+where $\mathbb{I}$ is the indicator function, $f(\sigma_{1},\sigma_{2},\rho)$ is a chosen criterion and $C$ is a threshold chosen as a hyperparameter.
+However, this methodology has notable limitations when applied to the reconstruction of geometrically complex static structures.
+In such cases, even static objects may produce substantial residuals, leading to their misclassification as dynamic elements and their subsequent masking.
+This misclassification ultimately degrades the reconstruction quality of static scene components.
+To address these limitations, we introduce necessary regularization constraints.
+We note that, even though we train our RUP only on a reference images, we can also obtain a per pixel uncertainty prediction $\hat{\Sigma}$ using an image reconstructed by Gaussian Splatting model.
+Because our model relies on semantic information of DINOv2 features, we should expect it to make a similar prediction in the static areas and a different one in areas corresponding to the dynamic objects.
+To estimate this discrepancy, we calculate the Kullback-Leibler (KL) divergence $D_{KL}(\mathcal{N}(0,\Sigma)||\mathcal{N}(0,\hat{\Sigma}))$, which takes following form for two normal distributions:
+$D_{KL}(\mathcal{N}(0,\Sigma)||\mathcal{N}(0,\hat{\Sigma}))=\frac{1}{2}(tr(\hat{\Sigma}^{-1}\Sigma)-ln(\frac{|\Sigma|}{|\hat{\Sigma}|})-2).$ (9)
 Fig.
-3 demonstrates the effectiveness of this approach: after 7000 iterations, the model with consistency loss produces precise masks for transient objects in the input image while generating minimal masks for the rendered image.
-In contrast, without the consistency term, the TMP exhibits numerous false positives in both input and rendered images, often misclassifying poorly reconstructed static regions as transient objects.
-This behavior validates that the consistency loss effectively distinguishes between genuine dynamic objects and reconstruction artifacts based on their semantic representations.
-Figure 3: The comparison of the TMP predictions with and without $\mathcal{L}_{\text{consist}}$ term.
-The quality of the predicted transient masks for the input image with consistency loss term is significantly better than without this term.
-The predicted masks by TMP are shown in green color.
+3 illustrates how this approach reduces false classifications in static regions.
+Unlike previous methods that obtain masks by estimating uncertainty, we instead utilize divergence.
+This allows us to incorporate additional information to enhance the consistency of our masks.
+Hence, our binary masks are obtained based on the new criterion:
+$M=\mathbb{I}(D_{KL}>C).$ (10)
+Figure 3: During the initial stages of reconstruction, RUP predicts high uncertainty in challenging regions such as backgrounds or high frequency details.
+However, since RUP relies exclusively on semantic information, calculating the divergence between reference uncertainty $\Sigma$ and reconstructed uncertainty $\hat{\Sigma}$ effectively suppresses these artifacts.
+Areas with divergence values above the threshold are highlighted in red, while the final predicted transient mask by RUP is shown in green.
+Notably, formula (9) is not exact since we are working with a folded multivariate distribution.
+We can add an ad hoc assumption that we observe only the absolute value of the residuals and its sign is random.
+Nevertheless, it has proven to be a useful heuristic.
+We leave a more careful derivation for future work.
 3.2.5 Training Stability
-To ensure stable training, we implement several key strategies.
-First, we detach $\hat{P}$ from the computational graph during consistency loss computation to prevent the model from learning to distinguish between ground truth and rendered images.
-During 3DGS optimization, there are periods when renders may be unreliable: at the beginning of training and after each opacity reset.
-Following  [15], we address this by implementing two key strategies.
-First, we delay the start of TMP training until the 3DGS optimization has completed its first 500 iterations, ensuring the initial scene reconstruction has reached sufficient quality.
-Second, after each opacity reset, we temporarily pause the TMP optimization for 250 iterations while keep training 3DGS, allowing the reconstruction to stabilize before resuming transient detection.
-Prior to bilinear interpolation, we dilate the masks by one pixel.
-This dilation step serves two purposes: it compensates for the coarse feature resolution of the foundation model (which processes images in patches) and helps address imprecise boundary predictions.
-These modifications ensure robust training and accurate transient object detection across diverse dynamic scenes.
+During 3DGS optimization, there are periods when renders may be unreliable, particularly at the beginning of training and after each opacity reset.
+Building on  [12], we address this by implementing two key strategies.
+First, we delay the start of RUP training until the 3DGS optimization has completed its first 500 iterations, ensuring the initial scene reconstruction has reached sufficient quality.
+Second, after each opacity reset, we temporarily pause the RUP optimization for 250 iterations while keep training 3DGS, allowing the reconstruction to stabilize before resuming transient detection.
+We also use scheduled sampling technique from SpotLessSplats [22].
+3.2.6 Mask Dilation
+We also dilate our masks, depending on the resolution of the scene.
+This dilation step serves multiple purposes, primarily covering shadows and reflections caused by objects.
+These modifications ensure robust training and accurate detection of transient objects in diverse dynamic scenes.
 3.3 Transient Mask Refinement
-Our transient area detection pipeline is robust for current benchmarks, where transient object is always dynamic and changes it’s position from frame to frame.
-But for semi-transient objects, which may do not change position for some frames it fails and masks only parts of video when it’s dynamic.
-To address these issue, we introduce a masks propagation process that refines transient masks into temporally consistent, accurate masks with high-resolution boundaries across the entire video sequence through refinement and propagation.
-Each segmentation consists of a binary mask defining the object’s spatial extent and a unique label that remains consistent throughout the video sequence.
+Our transient area detection pipeline is robust for current benchmarks, where transient objects are always dynamic and change their positions from frame to frame.
+However, for semi-transient objects, which may not change positions for some frames, it fails and masks only parts of the video when they are dynamic.
+To address this issue, we introduce a mask propagation process that refines transient masks into temporally consistent, accurate masks with high-resolution boundaries across the entire video sequence through refinement and propagation.
+Each segmentation consists of a binary mask that defines the object’s spatial extent and a unique label, consistent throughout the video sequence.
 3.3.1 Spatial Refinement
-We employ the Segment Anything Model (SAM) [14] to refine our transient maps $P_{i}$ into more precise masks $M^{\prime}_{i}$.
+We use the Segment Anything Model (SAM) [11] to refine our transient maps, $P_{i}$, into more precise masks, $M^{\prime}_{i}$.
 For each connected component $C_{i}^{k}$ in $P_{i}$, we sample up to ten points as prompts for SAM, leveraging its ability to generate high-quality segmentations from sparse inputs to extract a set of object-aware masks $M^{\prime}_{i}=\{M^{\prime}{{}_{i}^{j}}\}_{j=1}^{L_{i}}$, where $L_{i}$ is the number of predicted masks for image $I_{i}$.
 Due to potential inaccuracies in the boundaries of our masks, some sampled points might occasionally fall on the background rather than the object itself (e.g., a point sampled between the legs of a person).
-To address this issue, we filter the predicted masks based on their local coverage score:
-$CS_{\text{local},i}=\frac{|P_{i}\cap M^{\prime}{{}_{i}^{j}}|}{|M^{\prime}{{}_{i}^{j}}|}$ (4)
+To address this, we filter the predicted masks based on their local coverage score:
+$CS_{\text{local},i}=\frac{|P_{i}\cap M^{\prime}{{}_{i}^{j}}|}{|M^{\prime}{{}_{i}^{j}}|}.$ (11)
 We keep only masks that satisfy $CS_{\text{local},i}>\lambda_{\text{cov}}^{\text{ref}}$, forming the refined set $M^{\prime}_{i}=\{M^{\prime}{{}_{i}^{j}}|CS_{\text{local},i}>\lambda_{\text{cov}}^{\text{ref}}\}$.
 3.3.2 Temporal Refinement
-To address potential false negatives, we propagate the refined masks $\{M^{\prime}_{i}\}_{i=1}^{N}$ throughout the video using SAM2 [29] to obtain more consistent masks $\{M_{i}\}_{i=1}^{N}$.
+To address potential false negatives, we propagate the refined masks, $\{M^{\prime}_{i}\}_{i=1}^{N}$, throughout the video using SAM2 [19] to obtain more consistent masks, $\{M_{i}\}_{i=1}^{N}$.
 Our propagation process consists of three stages:
   1. 1.
-    Forward Propagation: Iterating from the first to the last frame, to propagate segmentation masks forward.
+    Forward Propagation: Iterating from the first frame to the last, propagating the segmentation masks forward.
   2. 2.
-    Backward Propagation: Iterating Last-to-first frame, to propagate information from future frames backwards.
+    Backward Propagation: Iterating from the last frame to the first, propagating information from future frames backward.
   3. 3.
-    Final Propagation: Final Propagation: A final first-to-last pass, considering both past and future frames as context, which helps resolve any temporal inconsistencies.
-To manage computational resources efficiently, we introduce a memory size parameter $N_{m}$, which limits the number of frames considered during propagation.
-At each step, we only maintain and use the segmentations from the $N_{m}$ most nearest frames, balancing between temporal consistency and memory constraints.
-Throughout propagation, we manage mask intersections to ensure coherent segmentation.
-For any pair of masks $M_{i}^{l}$ and $M_{i}^{m}$ with $IoU(M_{i}^{l},M_{i}^{m})>\lambda_{merge}$, we merge them into a single mask, assigning the lower of the two original labels to maintain consistency.
+    Final Propagation: A final first-to-last pass, considering both past and future frames as context, which helps to resolve temporal inconsistencies.
+To manage computational resources efficiently, we introduce a memory size parameter, $N_{m}$, which limits the number of frames considered during propagation.
+At each step, we maintain and use segmentations from $N_{m}$ nearest frames, balancing temporal consistency with memory constraints.
+During propagation, we manage mask intersections to ensure consistent segmentation.
+For any pair of masks $M_{i}^{l}$ and $M_{i}^{m}$ where $IoU(M_{i}^{l},M_{i}^{m})>\lambda_{merge}$, we merge them into a single mask, assigning the lower of the two original labels to maintain consistency.
 3.3.3 Dynamic Object Filtration
-To filter out false positive transients and ensure robust detection, we introduce a Stability Ratio (SR) metric that combines both spatial overlap accuracy and temporal consistency.
+To filter out false positive transients and ensure robust detection, we introduce the Stability Ratio (SR) metric, which combines spatial overlap accuracy and temporal consistency.
 For each detected object, the SR is calculated as $SR=\frac{1}{N}\sum_{i=1}^{N}(R_{i}\cdot CS_{\text{global},i})$, where $N$ is the number of valid frames, $R_{i}$ is the mean value of the absolute difference between ground truth and rendered images within the masked region in frame $i$, and $CS_{\text{global},i}=|P_{i}\cap M_{i}|/|M_{\text{max}}|$ is the global coverage score.
 Here, $P_{i}$ represents the prompt mask in frame $i$, $M_{i}$ is the segmentation mask, and $M_{\text{max}}$ is the maximum size of the object mask across all frames.
-This global score helps evaluate the object’s consistency relative to its largest observed size.
+This global score evaluates the object’s consistency relative to its largest observed size.
 A frame is considered valid and contributes to the SR calculation only if its local coverage score (Eq.
-4) exceeds the validation threshold $\lambda_{\text{cov}}^{\text{val}}$.
+11) exceeds the validation threshold $\lambda_{\text{cov}}^{\text{val}}$.
 Objects with $SR$ below a threshold $\lambda_{SR}$ are filtered out as potential false detections.
@@
 3DGS tends to generate floating artifacts (”floaters”) near the camera, particularly in challenging regions like those identified by transient masks.
-These artifacts can saturate gradients and degrade overall reconstruction quality.
+These artifacts can saturate gradients, thereby degrading overall reconstruction quality.
 We address this issue through depth-aware regularization.
-We render depth D for each pixel using alpha compositing similar to color rendering: $D=\sum_{i=1}^{M}T_{i}\alpha_{i}d_{i}$, where $d_{i}$ is the depth value of the i-th Gaussian, $T_{i}$ is the accumulated transmittance, and $\alpha_{i}$ is the opacity value.
-To suppress floating artifacts while preserving sharp depth discontinuities at object boundaries, we apply anisotropic total variation (TV) regularization to the rendered depth maps: $\mathcal{L}_{\text{depth}}=\text{mean}(|\nabla_{x}D|)+\text{mean}(|\nabla_{y}D|)$, where p indexes over pixels, $\nabla_{x}$ and $\nabla_{y}$ are spatial gradients in x and y directions respectively.
+We render the depth $D$ for each pixel using alpha compositing, similar to color rendering: $D=\sum_{i=1}^{M}T_{i}\alpha_{i}d_{i}$, where $d_{i}$ is the depth value of the i-th Gaussian, $T_{i}$ is the accumulated transmittance, and $\alpha_{i}$ is the opacity value.
+To suppress floating artifacts while preserving sharp depth discontinuities at object boundaries, we apply anisotropic total variation (TV) regularization to the rendered depth map: $\mathcal{L}_{\text{depth}}=\text{mean}(|\nabla_{x}D|)+\text{mean}(|\nabla_{y}D|)$, where $\nabla_{x}$ and $\nabla_{y}$ are spatial gradients in $x$ and $y$ directions respectively.
 3.5 Masked Gaussian Splatting Optimization
 The final step involves training the Gaussian Splatting model with the obtained masks $\{M_{i}\}_{i=1}^{N}$ for transients. Let $M_{i}$ be the binary mask for frame $i$, defined as:
-$M_{i}(x,y)=\begin{cases}1&\text{if (x,y) is in an occluded area}\\ 0&\text{if (x,y) is in a static area},\end{cases}$ (5)
+$M_{i}(x,y)=\begin{cases}1&\text{if $(x,y)$ is in an occluded area},\\ 0&\text{if $(x,y)$ is in a static area},\end{cases}$ (12)
 where $(x,y)$ represents pixel coordinates in the image.
 We apply binary dilation to $M_{i}$ for $N_{e}$ iterations, yielding $M^{*}_{i}$.
-This operation creates a buffer zone around detected dynamic objects, enhancing the robustness of our static scene reconstruction.
+This operation creates a buffer zone around detected dynamic objects, improving the robustness of our static scene reconstruction.
 The final loss for 3DGS is:
 $\begin{split}\mathcal{L}_{\text{masked}}=\lambda_{\text{SSIM}}\cdot L_{\text{SSIM}}(I_{i}\odot\overline{M}^{*}_{i},\hat{I_{i}}\odot\overline{M}^{*}_{i})+\\ \lambda_{\text{L1}}\cdot\left|\left|\overline{M}^{*}_{i}\odot(I_{i}-\hat{I_{i}})\right|\right|_{1}+\lambda_{\text{depth}}\mathcal{L}_{\text{depth}},\end{split}$ (13)
-where $I_{i}$, $\hat{I_{i}}$, $\mathcal{L}_{depth}$, $\odot$ and $\|\cdot\|_{1}$ are the same as were defined previously, $L_{SSIM}$ is a structural similarity loss, $\overline{M}^{*}_{i}$ is a negation of $M^{*}_{i}$ that represents static background and $\lambda_{\text{SSIM}}$, $\lambda_{\text{L1}}$ and $\lambda_{\text{depth}}$ are weighting factors.
+where $I_{i}$, $\hat{I_{i}}$ are reference images and their reconstructions, $\odot$ is the Hadamard product, $\|\cdot\|_{1}$ is L1 norm, $L_{SSIM}$ is a structural similarity loss, $\overline{M}^{*}_{i}$ is a negation of $M^{*}_{i}$ that represents a static background and $\lambda_{\text{SSIM}}$, $\lambda_{\text{L1}}$ and $\lambda_{\text{depth}}$ are weighting factors.
 This formulation allows the model to focus on static scene elements, effectively handling dynamic objects in the reconstruction process.
-By integrating these steps, our method robustly reconstructs static scenes from unconstrained videos, effectively handling transient distractors.
+By integrating these steps, our method reconstructs static scenes robustly from unconstrained videos while effectively handling transient distractors.
 4 Experiments
+Figure 4: Qualitative results on the On-the-go dataset. Our method outperforms existing approaches in detecting transient objects. Predicted transient masks are shown in green.
+Figure 5: Qualitative results on the T-3DGS dataset. Our method produces cleaner transient masks and further refines them using the (TMR) module.
 Mountain Fountain Corner Patio Spot Patio High Mean
-PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
-NeRF On-the-go [31] 20.15 0.64 20.11 0.61 24.22 0.81 20.78 0.75 23.33 0.79 21.41 0.72 21.67 0.72
-3DGS [12] 20.09 0.70 20.11 0.69 23.51 0.85 15.87 0.68 20.78 0.80 17.80 0.69 19.69 0.74
-Robust3DGS [43] 16.97 0.61 18.18 0.59 23.47 0.85 21.33 0.85 22.61 0.88 21.81 0.82 20.73 0.77
-WildGaussians [15] 20.77 0.70 20.74 0.67 25.79 0.88 21.77 0.85 24.39 0.88 22.36 0.80 22.64 0.80
-SpotLessSplats [34] 21.25 0.66 20.49 0.63 25.59 0.85 21.13 0.80 24.13 0.78 22.18 0.76 22.46 0.75
-Ours 20.89 0.72 21.01 0.69 26.64 0.90 21.55 0.86 25.96 0.91 23.18 0.84 23.21 0.82
-Table 1: Quantitative comparison on On-the-go dataset [31].
+PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+NeRF On-the-go [20] 20.15 0.64 0.26 20.11 0.61 0.31 24.22 0.81 0.19 20.78 0.75 0.22 23.33 0.79 0.19 21.41 0.72 0.24 21.67 0.72 0.24
+3DGS [9] 19.40 0.66 0.21 19.96 0.66 0.19 20.90 0.71 0.24 17.48 0.70 0.20 20.77 0.69 0.32 17.29 0.60 0.36 19.30 0.67 0.25
+Robust3DGS [28] 16.97 0.61 0.31 18.18 0.59 0.32 23.47 0.85 0.10 21.33 0.85 0.07 22.61 0.88 0.12 21.81 0.82 0.17 20.73 0.77 0.19
+WildGaussians [12] 20.77 0.70 0.23 20.74 0.67 0.21 25.79 0.88 0.09 21.77 0.85 0.07 24.39 0.88 0.10 22.36 0.80 0.17 22.64 0.80 0.15
+SpotLessSplats [22] 21.25 0.66 0.24 20.49 0.63 0.24 25.59 0.85 0.12 21.13 0.80 0.08 24.13 0.78 0.18 22.18 0.76 0.20 22.46 0.75 0.18
+Ours 21.11 0.71 0.22 20.94 0.69 0.21 26.46 0.90 0.12 21.95 0.87 0.10 25.78 0.90 0.12 22.76 0.83 0.17 23.17 0.82 0.16
+Table 1: Quantitative comparison on the On-the-go dataset [20].
 Lab 1 Lab 2 Library Anti-Stress Office Mean
-PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
-3DGS [12] 24.49 0.91 20.42 0.87 20.08 0.89 20.45 0.86 26.96 0.94 22.48 0.89
-Robust3DGS [43] 25.35 0.93 24.74 0.92 24.33 0.93 22.95 0.91 28.52 0.96 25.18 0.93
-WildGaussians [15] 25.71 0.92 23.68 0.91 24.65 0.92 21.69 0.89 28.89 0.95 24.92 0.92
-SpotLessSplats [34] 25.28 0.91 24.63 0.90 24.11 0.91 22.22 0.90 28.08 0.92 24.86 0.91
-Ours w/o TMR 25.69 0.93 24.74 0.92 23.82 0.92 24.21 0.92 29.01 0.95 25.49 0.93
-Ours w/ TMR 27.71 0.95 25.42 0.93 28.34 0.97 28.79 0.97 29.87 0.96 28.03 0.96
-Table 2: Quantitative comparison on our T-3DGS dataset.
+PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+3DGS [9] 24.49 0.91 - 20.42 0.87 - 20.08 0.89 - 20.45 0.86 - 26.96 0.94 - 22.48 0.89 -
+Robust3DGS [28] 25.35 0.93 0.09 24.74 0.92 0.10 24.33 0.93 0.08 22.95 0.91 0.10 28.52 0.96 0.06 25.18 0.93 0.09
+WildGaussians [12] 25.71 0.92 0.08 23.68 0.91 0.11 24.65 0.92 0.09 21.69 0.89 0.09 28.89 0.95 0.04 24.92 0.92 0.08
+SpotLessSplats [22] 25.28 0.91 0.08 24.63 0.90 0.09 24.11 0.91 0.09 22.22 0.90 0.10 28.08 0.92 0.05 24.86 0.91 0.08
+Ours w/o TMR 25.77 0.93 0.09 24.67 0.92 0.10 24.67 0.93 0.08 24.07 0.92 0.09 29.36 0.95 0.05 25.71 0.93 0.08
+Ours w/ TMR 27.76 0.96 0.02 25.54 0.93 0.06 28.25 0.97 0.02 29.01 0.96 0.02 29.85 0.96 0.02 28.08 0.96 0.03
+Table 2: Quantitative comparison on the T-3DGS dataset.
 We evaluate our proposed T-3DGS model on various datasets captured in uncontrolled settings and filled with diverse distractors.
-We perform qualitative and quantitative comparisons of our method with other SOTA methods.
-Finally, we provide an ablation study of our design and loss choices and discuss the limitations of the proposed method.
+We perform qualitative and quantitative comparisons against state-of-the-art methods.
+Finally, we provide an ablation study of architectural and loss function choices.
+We discuss the limitations of the proposed method in the Supplementary Material.
 Datasets.
 We evaluate our model on three challenging datasets.
-The NeRF On-the-go dataset [31] contains 4 outdoor and 2 indoor sparsely captured scenes with different levels of occlusions (from 5% to over 30%) and minimal changes in appearance.
-The RobustNeRF dataset [33] contains 5 indoor scenes with unintentional changes during the capture process.
-Those changes include both transient objects that appear and disappear without forming a coherent video sequence (i.e., do not have a clear temporal ordering) and dynamic objects (e.g., floating balloons).
+The NeRF On-the-go dataset [20] contains four outdoor and two indoor sparsely captured scenes with different levels of occlusion (from 5% to over 30%) and minimal appearance changes.
+The RobustNeRF dataset [21] contains five indoor scenes with unintentional changes during the capture process.
+These changes include transient objects that appear and disappear without a consistent temporal order, as well as dynamic objects (e.g., floating balloons).
 Additionally, we introduce our novel T-3DGS dataset.
@@
 Generally, dynamic objects in our videos are walking people and various small objects.
-However, unlike previous datasets, all scenes include various adversarial cases, such as transient, semi-transient, and slow-moving objects.
+However, unlike previous datasets, all scenes incorporate challenging cases, including transient, semi-transient, and slow-moving objects.
 Baselines.
-We compare our model against vanilla 3D Gaussian Splatting [12] and against the current state-of-the-art method SpotLessSplats [34].
+We compare our model against vanilla 3D Gaussian Splatting [9] and the current state-of-the-art method, SpotLessSplats [22].
 We further include WildGaussians [12] and Robust3DGS [28] as baselines.
-To compare different models we adopt commonly used PSNR and SSIM [45] metrics for evaluation.
+To compare different models, we use commonly used PSNR, SSIM [30] and LPIPS metrics for evaluation.
 Implementation details.
-All our experiments are conducted following the training setup from the official 3DGS implementation.
-We train our models for 30K iterations, using the Adam optimizer with a learning rate of 1e-3 for the TMP.
+All our experiments are conducted in accordance with the training setup from the official 3DGS implementation.
+We train our models for 30K iterations, using the Adam optimizer with a learning rate of 1e-3 for the RUP.
 The depth regularization loss $\mathcal{L}_{\text{depth}}$ is activated after the first 500 iterations, allowing the 3DGS to establish initial geometry reconstruction.
-For the experiments with mask propagation, we first train the TMP for 7000 iterations.
-At that point we pause the training to propagate the transient masks.
+For the experiments with mask propagation, we first train the RUP for 7000 iterations.
+At that point, we pause the training to propagate the transient masks.
 Subsequently, we initiate a new training procedure using the propagated masks, keeping all other parameters the same as the original training setup.
-Thanks to our consistency term, we achieve robust results with a fixed regularization parameter $\lambda_{prior}=0.1$ across all scenes, eliminating the need for per-scene parameter tuning that was required in previous approaches.
-Figure 4: Qualitative results on the On-the-go dataset. Our method detects transient objects better than our competitors. Predicted masks are shown in green.
-Figure 5: Qualitative results on T-3DGS dataset. Our method detects cleaner transient masks and improves them further by mask propagation TMR module. Predicted masks are shown in green.
+We dilate all our masks by 10 pixels, except for the Patio scene, where we use the original mask due to its low resolution.
 4.1 Quantitative Comparisons
@@
 We report results on On-the-go and T-3DGS datasets in Tab.
-1 and 2, respectively, and we move evaluation results on RobustNeRF dataset to Supplementary Material as it is the least challenging.
+1 and 2, respectively, and we move the evaluation results of RobustNeRF dataset to the Supplementary Material as it presents the least challenge.
 As shown in Tab.
-1 and 2 our method generally outperforms current SOTA methods, in particular, our method is robust to changes in distant and high-frequency details.
+1 and 2, our method generally outperforms current SOTA methods.
+In particular, our method is robust to changes in distant and high-frequency details.
 In Tab.
-1 we run our method directly on masks predicted by TMP module without mask propagator.
+1 we run our method directly on masks predicted by RUP module without mask propagator.
 While current SOTA methods struggle to detect semi-transient objects (Tab.
-2), our proposed transient network TMP achieves higher performance minimizing false predictions.
-The integration of the SAM-based mask propagation TMR module further enhances our results in scene reconstruction filled with semi-transient objects, providing more accurate and reliable reconstructions.
+2), our proposed transient network RUP achieves higher performance by minimizing false predictions.
+The integration of the SAM-based mask propagation TMR module further enhances our results in scenes containing semi-transient objects, providing more accurate and reliable reconstructions.
 4.2 Qualitative Comparisons
-For qualitative comparison, we compare our method to SpotLessSplats [34] and Robust3DGS [43].
+For qualitative evaluation, we compare our method to SpotLessSplats [22], Robust3DGS [28], and WildGaussians [12].
 Fig.
-4 and 5 show that our method has very few false negatives and robustly detects transients.
-For example, on the On-the-go dataset most of the methods struggle with high-frequency details and distant objects, as both generally reconstruct slower than the rest of the scene, and therefore confuse RGB residuals-based and similar methods.
-However, due to the introduction of our robust loss, such artifacts generally disappear from our dynamic maps.
-While our masks are less sharp, as we upsample them from low-resolution DINOv2 patches, techniques such as guided upsampling can improve performance of our method, however we leave this topic to following works.
-It should be noted, that SpotLessSplats [34] use features obtained from higher-resolution images, while we compute our features on a lower resolution, the same resolution used for training 3DGS.
-For our T-3DGS dataset, we additionally utilize the SAM-based mask propagation module to propagate object-aware masks for semi-transient objects as shown in Fig.
+4 and 5 demonstrate that our method minimizes false negatives and effectively detects transients.
+For example, in the On-the-go dataset, most methods struggle with high-frequency details and distant objects, as these elements are typically reconstructed more slowly than the rest of the scene, leading to inaccuracies in RGB residual-based approaches.
+However, due to our robust loss function, such artifacts are largely eliminated from our dynamic maps.
+Notably, SpotLessSplats uses features obtained from higher-resolution images, while we extract features at a lower resolution, the same resolution used for training 3DGS.
+For our T-3DGS dataset, we additionally utilize the SAM-based mask propagation module to propagate object-aware masks for semi-transient objects, as shown in Fig.
 5.
-While most methods would benefit from such a technique in principle, our masks have higher quality and fewer wrong detections.
-Applying mask propagation to other methods risks propagating errors.
+Although most methods would theoretically benefit from this technique, our masks are of higher quality and result in fewer incorrect detections.
+Applying mask propagation to other methods may introduce error propagation, as demonstrated in the Supplementary Material.
 On-the-go dataset
 PSNR $\uparrow$ SSIM $\uparrow$
-w/o dilation 22.65 0.82
-w/o $\mathcal{L}_{\text{consist}}$ 16.76 0.69
-w/o $\mathcal{L}_{\text{depth}}$ 23.02 0.82
-Ours (full) 23.21 0.82
-Table 3: We ablate the importance of each component of our method on On-the-go dataset. We report the mean over all scenes.
+GT masks w/o $\mathcal{L}_{\text{depth}}$ 22.84 0.82
+GT masks w/ $\mathcal{L}_{\text{depth}}$ and dilation 23.43 0.81
+Ours w/o dilation and $\mathcal{L}_{\text{depth}}$ 22.60 0.80
+Ours w/o dilation 22.88 0.80
+Ours (full) 23.41 0.81
+Table 3: We evaluate the importance of each component of our method on the On-the-go dataset. We report the average performance across all scenes.
 4.3 Ablation Study
-We show the importance of each component in our T-3DGS model, separately disabling each of them.
-We report ablation results in Tab.
-3 on On-the-go dataset.
-In the first experiment, we show the importance of the dilation before upsampling per-feature probability masks $P^{f}$.
-Without dilation the produced masks keep shrinking during the optimization process.
-This leads to a quality reduction in the renderings.
-In the second experiment, we turn off consistency loss term $\mathcal{L}_{\text{consist}}$.
-Our results show that adding this loss term improves the quality of renderings across all datasets.
-In our last experiment, we ablate the regularization term on depth $\mathcal{L}_{\text{depth}}$.
-Even though the metrics are only slightly improved, adding this term helps to remove floating artifacts in the reconstructed scene.
-We include visualizations to support our statements in Supplementary Material.
+We present ablation results in Table 3 for the On-the-go dataset, excluding the Patio scene due to its low resolution.
+We evaluate our method under the following conditions: (1) without mask dilation, (2) without mask dilation and $\mathcal{L}_{\text{depth}}$, and (3) with both components enabled.
+Additionally, we report results obtained with ground truth masks while separately disabling $\mathcal{L}_{\text{depth}}$ and mask dilation by 10 pixels.
+Even when using ground truth masks, dilation noticeably enhances performance.
+This contradicts the assumptions made by NeRF-HuGS [2] and Robust3DGS [28], as exact masks do not yield optimal performance metrics.
+Furthermore, mask dilation aids RUP training by ensuring that all transient objects are fully covered.
+We also note that our results align very closely with those obtained using GT masks, suggesting that more challenging datasets are required.
 5 Conclusion
-In this work, we have presented novel T-3DGS method for 3D scene reconstruction using Gaussian Splatting by effectively removing foreground dynamic distractors from input videos.
-The key to the success are unsupervised classification network to identify transient distractors through novel consistency loss term and the mask propagation approach that improves the boundary and temporal consistency of the predicted transient masks.
-Our results show that our method achieves significant improvement over the state-of-the-art methods on densely captured videos populated with semi-transient distractors.
-We believe our method is a valuable contribution towards widespread usage of 3DGS for robust 3D scene reconstruction from real-world videos captured in uncontrolled settings.
-[…]
+In this work, we have presented the novel T-3DGS method for 3D scene reconstruction using Gaussian Splatting by effectively filtering out foreground dynamic distractors from input videos.
+By integrating an unsupervised classification network with bivariate uncertainty modeling, KL divergence regularization, and a mask propagation strategy, our method achieves superior temporal coherence and boundary accuracy.
+Evaluations on both sparsely and densely captured datasets confirm significant improvements over state-of-the-art approaches.
+We believe our method represents a significant step toward the broader adoption of 3DGS for robust 3D scene reconstruction from real-world videos captured in uncontrolled settings.
 Supplementary Material
-A Additional Implementation Details
-For mask filtering and refinement in Sec.
-3.3.1, we set $\lambda_{\text{cov}}^{\text{ref}}=0.7$ for initial mask refinement and $\lambda_{\text{cov}}^{\text{val}}=0.7$ for validation during object filtration.
+Appendix A Limitations
+We use features upscaled by FeatUP [6] to compute cosine distance, and while it is better than simple bilinear interpolation, it is relatively slow and gives fairly noisy results.
+Utilizing alternative ways to measure per pixel errors might improve both speed and accuracy of the method.
+Additionally, the temporal refinement process is constrained by a memory window of $N_{m}$ frames, which means that if an object disappears for more than $N_{m}$ frames and then reappears, it will be treated as a new instance with a different label.
+This can lead to inconsistent tracking and potentially affect the filtering process, especially for semi-transient objects that may temporarily leave the scene.
+Furthermore, our current filtering approach using global coverage scores may incorrectly filter out valid dynamic objects that undergo significant size changes, such as objects moving towards or away from the camera, or those experiencing perspective changes.
+We leave it as a future work.
+Appendix B Additional Implementation Details
+In Sec.
+3.3.1, for mask filtering and refinement, we set $\lambda_{\text{cov}}^{\text{ref}}=0.7$ for initial mask refinement and $\lambda_{\text{cov}}^{\text{val}}=0.7$ for validation during object filtration.
 For temporal refinement in Sec.
@@
 For the final mask dilation step, we perform $N_{e}=5$ iterations of binary dilation.
-Also, the mask merging threshold $\lambda_{merge}$ is set to 0.9, and the stability ratio threshold $\lambda_{SR}$ to 0.08 in Sec.
+In addition, the mask merging threshold $\lambda_{merge}$ is set to 0.9, and the stability ratio threshold $\lambda_{SR}$ to 0.08 in Sec.
 3.3.3.
-B Evaluation on RobustNeRF Dataset
-We provide evaluation on the RobustNeRF dataset [33].
+Our model consists of repeating blocks.
+We fist use bilinear interpolation to increase the resolution of our features by two.
+We then apply a simple 3 by 3 convolutional layer that also decreases feature size by a factor of two.
+We then apply layer normalization followed by the GELU non-linearity.
+We repeat this sequence three times.
+After that we project our features with 1 by 1 convolution to obtain logits.
+We use softplus for $\sigma_{1}$, $\sigma_{2}$ and tanh for $\rho$.
+The normalization layer is crucial for improving the numerical stability that arises due to matrix $\Sigma$ being potentially ill-conditioned.
+Appendix C Evaluation on RobustNeRF Dataset
+We evaluate our method on the RobustNeRF dataset [21].
 As shown in Tab.
-4 our method generally outperforms 3DGS [12], Robust 3DGS [43], WildGaussians [15], and shows similar performance comparing to SpotLessSplats [34].
-We run our method directly on masks predicted by TMP module without mask propagator (TMR).
-In general, the dataset does not seem to be challenging enough to highlight the differences between different methods.
+4 our method generally outperforms 3DGS [9], Robust 3DGS [28], WildGaussians [12], and shows similar performance compared to SpotLessSplats [22].
+We run our method directly on masks predicted by RUP module without mask propagator (TMR).
+Overall, the dataset does not appear to be sufficiently challenging to differentiate between the methods.
 Android Statue Crab (1) Crab (2) Yoda Mean
@@
 SpotLessSplats [22] 24.45 0.79 22.50 0.80 35.45 0.95 33.29 0.94 33.55 0.94 29.85 0.88
-Ours 25.04 0.84 22.90 0.87 35.81 0.95 33.36 0.93 32.09 0.92 29.84 0.90
-Table 4: Quantitative comparison on RobustNeRF dataset [33].
-Figure 6: Qualitative results on On-the-go dataset on the training frames. Our method produces higher-quality renderings without artifacts.
-Figure 7: Qualitative results on On-the-go dataset on the testing frames. Our method produces higher-quality renderings without artifacts.
-C More Qualitative Comparisons
-For qualitative comparison, we compare our method to SpotLessSplats [34] and Robust3DGS [43].
+Ours 25.10 0.84 22.90 0.87 34.25 0.95 33.85 0.93 32.45 0.93 29.71 0.90
+Table 4: Quantitative comparison on the RobustNeRF dataset [21].
+Anti-Stress Lab (1) Lab (2) Library Office Mean
+PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
+WildGaussians w/o TMR 21.69 0.89 25.71 0.92 23.68 0.91 24.65 0.92 28.89 0.95 24.92 0.92
+WildGaussians w/ TMR 24.07 0.92 24.65 0.92 24.84 0.93 28.32 0.97 29.75 0.96 26.33 0.94
+Ours w/ TMR 28.79 0.97 27.71 0.95 25.42 0.93 28.34 0.97 29.87 0.96 28.03 0.96
+Table 5: Evaluation of WildGaussians with TMR module on the Transient-3DGS dataset.
+Figure 6: Qualitative results on the On-the-go dataset using the training frames. Our method produces higher-quality renderings without artifacts.
+Figure 7: Qualitative results on the On-the-go dataset using the testing frames. Our method produces higher-quality renderings without artifacts.
+Appendix D Additional Experiments with TMR Module
+Even though our proposed TMR module leverages SAM2 to propagate the transient masks, we would like to emphasize that our method enables mask propagation spatially and temporally consistent, thereby providing more accurate and reliable reconstruction.
+Table 5 presents an evaluation of the reconstruction quality of WildGaussians with our TMR module.
+First, we obtained the transient masks using WildGaussians.
+Then, we propagate them through our TMR module.
+Finally, we reconstruct the scenes based on the transient masks obtained in the previous step.
+Our evaluation shows that our method produces higher-quality results for most scenes, with comparable performance in the remaining ones.
+Our method, with the TMR module, outperforms WildGaussians with the TMR module on Anti-Stress, Lab (1), Lab (2) scenes.
+The TMR module generally enhances the reconstruction quality of the original WildGaussians, but it is limited because of the false positive transient detections that come from WildGaussians itself.
+Furthermore, we note that the hyperparameters of our TMR module are highly dependent on the dataset rather than the model.
+That makes our TMR module robust across the different methods.
+Appendix E More Qualitative Comparisons
+For qualitative comparison, we evaluate our method against SpotLessSplats [22], Robust3DGS [28], and WildGaussians [12].
 We provide corresponding renderings for the masks shown in the main paper in Sec.
@@
 Currently, most methods can produce fairly good reconstructions and avoid significant artifacts, so generally, most methods produce fairly similar results (at least in the absence of semi-transient objects and other adversarial cases).
-However, it is notable that compared to other residual-based methods, we avoid misclassifying high-frequency details and similar objects.
-This improvement can be attributed to our consistency loss term.
-Figure 8: We ablate the importance of each component of our method on On-the-go dataset on Corner scene.
-We provide qualitative comparison for masks (row 1), rendered images (row 2), and rendered depths (row 3).
-Figure 9: We ablate the importance of each component of our method on On-the-go dataset on Spot scene.
-We provide qualitative comparison for masks (row 1), rendered images (row 2), and rendered depths (row 3).
-D Additional Ablation Study
-We provide visualizations that show the importance of each component in our proposed T-3DGS model on On-the-go dataset [31]. We show our ablation visualizations in Fig. 8 and 9.
-Figure 10: The comparison of predicted masks for full and reduced scenes.
-Figure 11: The comparison of predicted masks and scene reconstructions during movement of semi-transient objects for different frames.
-E Handling Semi-Transient Objects
-Semi-transient objects are have not been properly addressed in 3D scene reconstruction.
-Our method makes a large improvement over previous work and can handle fairly complex scenarios.
-We provide some details about data capture and the methodology we employed with respect to semi-transient objects.
-We also discuss why both our consistency term as well as mask propagation are important for handling semi-transient objects.
-Additionally, we discuss some limitations of our proposed method.
-Some scenes in our dataset include reduced and full versions - in the reduced scenes the camera operator goes from one end of the scene to the other, while for the full scenes they also retrace their path and return to their initial positions, while semi-transient objects kept moving.
-As seen in Fig.
-10, our proposed TMR module is crucial to produce good results even for the reduced scenes, the more challenging scenario.
-For full scenes, the additional data significantly improves mask predictions for all models.
-We conclude that handling semi-transient objects is an important challenge for in-the-wild video processing.
-In order to provide the most challenging datasets and to rigorously compare different methods - one must consider not only the kinds of motion of dynamic objects, but also their movement with respect to the camera.
+Notably, compared to other residual-based methods, we avoid misclassifying high-frequency details and similar objects.
+Figure 8: Comparison of predicted masks for full and reduced scenes.
+Figure 9: Comparison of predicted masks and scene reconstructions during the movement of semi-transient objects across different frames.
+Appendix F Handling Semi-Transient Objects
+Semi-transient objects have not been properly addressed in 3D scene reconstruction.
+Our method represents a significant improvement over previous work and can handle relatively complex scenarios.
+We provide details on the data capture process and the methodology employed for handling semi-transient objects.
+We also discuss the importance of both divergence estimation and mask propagation in handling semi-transient objects.
+Additionally, we discuss the limitations of our proposed method.
+Our dataset includes two versions of some scenes: reduced and full.
+In reduced scenes, the camera operator moves from one end of the scene to the other.
+In full scenes, however, the operator retraces this path back to the starting position while semi-transient objects continue to move.
+As illustrated in Fig.
+8, our proposed TMR module is essential for achieving good results in reduced scenes, which are particularly challenging.
+In full scenes, the additional frames lead to significantly improved mask predictions for all models because transient objects remain visible for longer periods.
+When fewer frames capture the scene, many methods mistakenly classify these transient objects as static.
+Overall, our findings highlight that effectively handling semi-transient objects is a major challenge in in-the-wild video processing.
+To develop the most challenging datasets and to rigorously compare different methods, it is important to consider not only the types of motion dynamic objects exhibit but also their movement relative to the camera.
 As mentioned above, some of the scenes include semi-transient objects occluding the static scene for prolonged periods of time while remaining mostly still.
 As this period of time increases, semi-transient objects can effectively become static.
-While this effect might seem irrelevant to the detection of dynamic objects, this is not the case.
+Although this effect might seem irrelevant to the detection of dynamic objects, this is not the case.
 As shown in the Fig.
-11, most methods mask static background as if it was masking the semi-transient object.
-Notably, our method avoids this problem, and this highlights the importance of using both the consistency term and mask propagation algorithm we introduced.
-While the consistency term alone may still misclassify semi-transient objects, we avoid false classifications of static objects as dynamic.
-Therefore, our consistency term is crucial in order for the mask propagation to avoid introducing additional errors.
+9, most methods mask the static background as if it were masking the semi-transient object.
+Notably, because WildGaussians relies heavily on semantic information, it can ”propagate” the masks.
+However, this happens too late into the training process while our method avoids this problem, and this highlights the importance of using both divergence estimation and mask propagation algorithm we have proposed.
+Moreover, we aim to minimize false classifications of static objects as dynamic.
+As discussed earlier, even WildGaussians produces an excessive number of misclassifications for TMR.
+Therefore, our method is crucial for mask propagation to avoid introducing additional errors.
 This is in stark contrast to the competing methods, which have a lot more false positives.
-Mask propagation would introduce additional errors and may not help improving the overall quality.
+Mask propagation could introduce additional errors and might not contribute to overall quality improvement.
 Our method reliably removes transient and semi-transient distractors and successfully reconstructs static artifact-free 3D scenes.
-However, we have noticed that predicted masks are inflated due to the low resolution of the extracted feature maps, our method can give inconsistent results for small objects, since DINOv2 features are computed on patches.
-These problems could be addressed by utilizing feature extractors with higher-resolution feature maps or guided upsampling.
-Additionally, the temporal refinement process is constrained by a memory window of $N_{m}$ frames, which means that if an object disappears for more than $N_{m}$ frames and then reappears, it will be treated as a new instance with a different label.
+However, we have observed that predicted masks tend to be inflated due to the low resolution of the extracted feature maps.
+Our method can also produce inconsistent results for small objects, as DINOv2 features are computed on patches.
+These problems could be addressed by using feature extractors with higher-resolution feature maps or guided upsampling.
+Additionally, the temporal refinement process is limited by a memory window of $N_{m}$ frames, which means that if an object disappears for more than $N_{m}$ frames and then reappears, it will be treated as a new instance with a different label.
 This can lead to inconsistent tracking and potentially affect the filtering process, especially for semi-transient objects that may temporarily leave the scene.
 Furthermore, our current filtering approach using global coverage scores may incorrectly filter out valid dynamic objects that undergo significant size changes, such as objects moving towards or away from the camera, or those experiencing perspective changes.
-We leave it as future work.
+We leave this aspect for future work.